### v3

Abstract
Two-time-scale optimization is a framework introduced in Zeng et al. (2024) that abstracts a range of policy evaluation and policy optimization problems in reinforcement learning (RL). Akin to bi-level optimization under a particular type of stochastic oracle, the two-time-scale optimization framework has an upper level objective whose gradient evaluation depends on the solution of a lower level problem, which is to find the root of a strongly monotone operator. In this work, we propose a new method for solving two-time-scale optimization that achieves significantly faster convergence than the prior arts. The key idea of our approach is to leverage an averaging step to improve the estimates of the operators in both lower and upper levels before using them to update the decision variables. These additional averaging steps eliminate the direct coupling between the main variables, enabling the accelerated performance of our algorithm. We characterize the finite-time convergence rates of the proposed algorithm under various conditions of the underlying objective function, including strong convexity, Polyak-Lojasiewicz condition, and general non-convexity. These rates significantly improve over the best-known complexity of the standard two-time-scale stochastic approximation algorithm. When applied to RL, we show how the proposed algorithm specializes to novel online sample-based methods that surpass or match the performance of the existing state of the art. Finally, we support our theoretical results with numerical simulations in RL.
Keywords: Bi-level optimization, accelerated stochastic optimization, reinforcement learning
1 Introduction
Inspired by the two-time-scale optimization framework for reinforcement learning (RL) introduced in Zeng et al. (2024), we study the following optimization problem
$\theta^{\star}=\textstyle\mathop{\rm argmin}_{\theta\in\mathbb{R}^{d}}~{}h(\theta),$ (1)
where $h:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a continuously differentiable (possibly nonconvex) function. The gradient of $h$ can only be accessed through a special stochastic oracle $F:\mathbb{R}^{d}\times\mathbb{R}^{r}\times{\cal X}\rightarrow\mathbb{R}$. The three arguments to $F$ are the decision variable $\theta\in\mathbb{R}^{d}$, an auxiliary variable $\omega\in\mathbb{R}^{r}$, and a random variable $X$ over the sample space ${\cal X}$. Given any $\theta$, $F(\theta,\omega,X)$ returns an unbiased estimate of the gradient $\nabla h(\theta)$ only under a single “correct” setting of $\omega$ when the random variable $X$ follows some distribution $\xi$
$\nabla_{\theta}h(\theta)$ $=\mathbb{E}_{X\sim\xi}[F(\theta,\omega^{\star}(\theta),X)],$ (2)
where $\omega^{\star}(\theta)\in\mathbb{R}^{r}$ is defined through another stochastic sampling operator $G:\mathbb{R}^{d}\times\mathbb{R}^{r}\times{\cal X}\rightarrow\mathbb{R}$, as the solution to the equation
$\mathbb{E}_{X\sim\xi}[G(\theta,\omega^{\star}(\theta),X]=0.$ (3)
In this work, the operator $G$ is assumed to be strongly monotone with respect to $\omega$, implying that $\omega^{\star}(\theta)$ exists and is unique for any $\theta$. Given any $\omega\neq\omega^{\star}(\theta)$, the stochastic gradient $F(\theta,\omega,X)$ may be arbitrarily biased. By the first-order optimality condition, it is obvious that solving (1) is equivalent to finding a pair of solutions $(\theta^{\star},\omega^{\star})$ that solve the following two coupled nonlinear equations
$\left\{\begin{array}[]{ll}\mathbb{E}_{X\sim\xi}[F(\theta^{\star},\omega^{\star},X)]=0,\\ \mathbb{E}_{X\sim\xi}[G(\theta^{\star},\omega^{\star},X)]=0.\end{array}\right.$ (6)
As detailed in Section 4, a wide range of RL methods, such as gradient-based temporal difference learning (Sutton et al., 2008, 2009) and actor-critic algorithms for policy optimization (Yang et al., 2019; Wu et al., 2020; Chen and Zhao, 2024), aim to solve problems that are special cases of (6). Thus, by developing and analyzing algorithms for the unified framework (6), we can apply them to various RL problems and immediately deduce performance guarantees, without having to tailor analysis for each specific problem. Hong et al. (2023); Zeng et al. (2024) pioneer the effort to construct such a unified framework and study solving (6) with the classic two-time-scale stochastic approximation (SA) algorithm, in which the iterates $\theta_{k},\omega_{k}$ are introduced to track $\theta^{\star},\omega^{\star}$ and updated in the following simple manner
[…]
where $X_{k}$ are independently drawn from $\xi$ in every iteration. The step sizes $\alpha_{k}$ and $\beta_{k}$ need to be chosen carefully according to the structure of $h$, usually one much smaller than the other, hence leading to the name “two-time-scale”. Despite the simplicity and popularity of this algorithm, its convergence rate is (at least as known so far) sub-optimal when compared to standard SGD under the most common oracle where we have access to $H(\theta,X)\in\mathbb{R}^{d}$ such that
$\mathbb{E}_{X\in\xi}[H(\theta,X)]=\nabla h(\theta).$ (8)
This is due to the complex coupling between $\theta_{k},$ $\omega_{k},$ and the noise $X_{k}$, Taking strongly convex function $h$ as an example, standard SGD with $k$ queries of the stochastic gradient $H$ is able to reduce the optimality error to ${\cal O}(1/k)$, while the two-time-scale SA algorithm can only achieve ${\cal O}(1/k^{2/3})$.
Remark 1
Under a Lipschitz/Holder continuous gradient assumption on $\omega^{\star}$, Shen and Chen (2022); Han et al. (2024) show that the standard two-time-scale algorithm achieves ${\cal O}(1/k)$ for strongly convex functions. We do not make this assumption in our work.
1.1 Main Contributions
First, our work proposes a new variant of the two-time-scale SA algorithm, presented in Algorithm 1, which for the first time achieves the optimal convergence rates when solving (6) under function $h$ satisfying various common structural properties, namely, strong convexity, Polyak-Lojasiewicz (PL) condition, and general non-convexity. We summarize the rates as well as the convergence metrics and choices of step sizes in Table 1. Our key idea behind the innovation is to leverage an averaging technique to estimate the unknown operators $F,G$ through their samples before updating the decision variables. In other words, we “denoise” the estimate of the stochastic operators $F,G$, which leads to more stable updates of $\theta_{k}$ and $\omega_{k}$. This seemingly simple modification effectively decouples the direct interaction between $\theta_{k}$ and $\omega_{k}$ and allows us to choose more aggressive step sizes, resulting in the accelerated convergence rates as well as simplified and elegant proofs.
Second, we show how the results derived for the general framework apply to policy evaluation and optimization problem in reinforcement learning. We discuss three specific applications, the first of which is temporal difference learning with gradient correction (TDC) under linear and non-linear function approximation. In the linear setting, the objective is strongly convex and Algorithm 1 specializes to a variant of the TDC algorithm guaranteed to converge with rate $\widetilde{{\cal O}}(1/k)$. In the non-linear setting, the objective is non-convex and our convergence rate is $\widetilde{{\cal O}}(1/k^{1/2})$. These results match the best-known rates of TDC in both cases. The second application is sample-based policy optimization in the linear-quadratic regulator, which has an objective function satisfying the PL condition. The state-of-the-art online actor-critic algorithm for this problem has a complexity of $\widetilde{{\cal O}}(1/k^{2/3})$ while our algorithm achieves the much faster rate $\widetilde{{\cal O}}(1/k)$. The final application is sample-based policy optimization in an entropy-regularized MDP, the objective of which also observes the PL condition. The results in Zeng et al. (2024) imply that the classic two-time-scale SA algorithm finds the globally optimal policy with rate $\widetilde{{\cal O}}(1/k^{2/3})$. Our algorithm in this context enjoys a finite-time rate of $\widetilde{{\cal O}}(1/k)$, again significantly elevating the state of the art.
Table 1: Summary of finite-time complexity under various structural assumptions on $h$.
Structural property Metric Complexity in this paper Order of step sizes Best known complexity of standard two-time-scale SA
Strong Convexity $\mathbb{E}[\|\theta_{k}-\theta^{\star}\|^{2}]$ ${\cal O}(k^{-1})$ $k^{-1}$ ${\cal O}(k^{-\frac{2}{3}})$
PL Condition $\mathbb{E}[h(\theta_{k})-h(\theta^{\star})]$ ${\cal O}(k^{-1})$ $k^{-1}$ ${\cal O}(k^{-\frac{2}{3}})$
Non-Convexity $\min_{t\leq k}\mathbb{E}[\|\nabla h(\theta_{t})\|^{2}]$ ${\cal O}(k^{-\frac{1}{2}})$ $k^{-\frac{1}{2}}$ ${\cal O}(k^{-\frac{2}{5}})$
1.2 Related Work
This paper is closely related to the existing works on two-time-scale stochastic approximation, stochastic bi-level and composite optimization, as well as those that analyze abstracted or specific RL algorithms. We discuss the latest advances in these subjects to give context to our contributions.
Two-time-scale stochastic approximation. The aim of two-time-scale SA, pioneered by works including Borkar (1997); Konda and Borkar (1999); Mokkadem and Pelletier (2006b), is to solve a system of equations in the form of (6). Existing works in this domain primarily focus on the settings where $F$ and $G$ are linear operators (Konda and Tsitsiklis, 2004; Dalal et al., 2018; Doan and Romberg, 2019; Dalal et al., 2020; Gupta et al., 2019) or strongly monotone operators (Mokkadem and Pelletier, 2006a; Doan, 2022; Shen and Chen, 2022). Our work can be seen as a generalization of this line of work to the case where $F$ is not a strongly monotone operator but the gradient map of a function $h$ exhibiting various structure.
Stochastic bi-level and composite optimization. Bi-level optimization (Colson et al., 2007; Chen et al., 2022a, 2023) studies solve an optimization program of the form
$\textstyle\min_{x}f_{1}(x,y^{\star}(x))\quad\text{subject to }y^{\star}(x)\in\mathop{\rm argmin}_{y}f_{2}(x,y),$ (9)
or equivalently (by the first-order condition), finding a pair of stationary points $(x^{\star},y^{\star})$ that observes
$\nabla_{x}f_{1}(x^{\star},y^{\star})=0,\quad\nabla_{y}f_{2}(x^{\star},y^{\star})=0.$
This is a special case of our objective (6) with $G$ being a gradient map. In RL problems $G$ usually abstracts the Bellman backup operator associated with policy evaluation, which is well-known to not be the gradient of any function. In this sense, our framework is more general and suitable for modeling applications in RL.
As another related subject, stochastic composite optimization (Ghadimi and Lan, 2012; Wang et al., 2017; Chen et al., 2021) studies problems structured as
$\min_{x}g_{1}(g_{2}(x)).$
At a first glance, this optimization objective reduces to (9) by choosing $g_{1}=f_{1}$, $g_{2}(x)=(x,y^{\star}(x))$, and $y^{\star}(x)$ to be the minimizer of $f_{2}(x,\cdot)$. Nevertheless, there is a key difference in the stochastic oracle. The assumption in stochastic composite optimization is that $g_{2}$ is continuous differentiable and that we can access an oracle that returns a direct stochastic estimate of $\nabla g_{2}$. This assumption is restrictive in most RL applications, in which there is no guarantee on the differentiability of $g_{2}$ and only indirect information of $g_{2}$ is accessible.
Finite-Time Analysis of RL algorithms. Last but not least, we note the connection of our paper to the vast volume of recent literature that present finite-time and finite-sample analysis for various RL algorithms including temporal difference learning (Wang et al., 2021; Haque et al., 2023), actor-critic algorithms for standard MDP (Wu et al., 2020; Olshevsky and Gharesifard, 2023; Chen and Zhao, 2024), and actor-critic algorithms for the linear-quadratic regulator (Yang et al., 2019; Ju et al., 2022; Zhou and Lu, 2023). As samples in RL are sequentially drawn from a Markov chain, many of these works drop the i.i.d. noise assumption and realistically assume that the samples $X_{k}$ are Markovian. Markovian samples can either be “time-invariant” in the sense that the stationary distribution of the Markov chain does not change over time (for example in policy evaluation or off-policy learning) or “controlled” if the samples are obtained from an Markov chain whose stationary distribution changes as the policy is updated (for example in actor-critic algorithms). To focus on the main contribution which is the acceleration due to stable estimates of operators $F$ and $G$, our work assumes i.i.d. noise for simplicity. We note that the techniques to handle “time-invariant” and “controlled” Markovian noise have been well-known since Srikant and Ying (2019) and Zou et al. (2019). Extending the analysis of Algorithm 1 to the Markovian noise setting can be done in a straightforward way and only makes the bounds worse by a $\log(k)$ factor. We discuss this extension in more details in Remark 2.
Outline of the paper. The rest of the paper is organized as follows. In Section 2 we present the proposed fast two-time-scale optimization algorithm. In Section 3 we study the finite-time convergence rates of the algorithm under various functional structure. Section 4 discusses how the optimization framework (1)-(3) applies to RL problems and how the proposed algorithm solves these problems faster than (or as fast as) the prior arts. Finally, numerical simulations verifying the superiority of the proposed algorithm are presented in Section 5.
2 Faster Two-Time-Scale Stochastic Gradient Method
We formally present the fast two-time-scale stochastic gradient method in Algorithm 1, in which we maintain and iteratively update four variables: $\theta_{k}$ and $\omega_{k}$ are used to track $\theta^{\star}$ and $\omega^{\star}(\theta^{\star})$, while $f_{k}$ and $g_{k}$ estimate the stochastic operators $F$ and $G$ at the current iterates $(\theta_{k},\omega_{k})$. An averaging step is carried out on the estimate of the stochastic operators. Note that Algorithm 1 reduces to the standard two-time-scale SA algorithm (7) if we set $\lambda_{k}=1$ for all $k$. With a large $\lambda_{k}$ the stochastic operator $F(\theta_{k},\omega_{k},X_{k})$ estimates the true gradient $\nabla h(\theta_{k})$ with high variance. By carefully choosing $\lambda_{k}$ to decay sufficiently fast, the algorithm updates the main variables with much more stable gradient estimates, which is the key driver behind the accelerated convergence rates.
Algorithm 1 Fast Two-Time-Scale Stochastic Gradient Method
1:  Initialization: decision variables $\theta_{0}\in\mathbb{R}^{d}$ and $\omega_{0}$, operator estimation variables $f_{0}\in\mathbb{R}^{d}$ and $g_{0}\in\mathbb{R}^{r}$, step size sequences $\{\alpha_{k},\beta_{k},\lambda_{k}\}$
2:  for $k=0,1,\cdots,K-1$ do
3:     Independent sample draw $X_{k}\sim\xi$
4:     Decision variable update:
[…]
5:     Gradient estimation variable update:
[…]
6:  end for
3 Finite-Time Complexity
In this section we study the finite-time complexity of Algorithm 1 under various structural properties of the objective function, namely, strong convexity, PL condition, and general non-convexity. Since the algorithm draws exactly one sample in each iteration, the time complexity translates to an equivalent sample complexity. We first state the main technical assumptions which are all fairly standard in the existing literature.
Assumption 1 (Lipschitz Continuity)
There exists a positive constant $L$ such that for any $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$, $\omega_{1},\omega_{2}\in\mathbb{R}^{r}$, and $X\in{\cal X}$
$\left\|F(\theta_{1},\omega_{1},X)-F(\theta_{2},\omega_{2},X)\right\|$ $\leq L\left(\|\theta_{1}-\theta_{2}\|+\|\omega_{1}-\omega_{2}\|\right),$
$\left\|G(\theta_{1},\omega_{1},X)-G(\theta_{2},\omega_{2},X)\right\|$ $\leq L\left(\|\theta_{1}-\theta_{2}\|+\|\omega_{1}-\omega_{2}\|\right),$
$\left\|\nabla h(\theta_{1})-\nabla h(\theta_{2})\right\|$ $\leq L\|\theta_{1}-\theta_{2}\|,$
$\|\omega^{\star}(\theta_{1})-\omega^{\star}(\theta_{2})\|$ $\leq L\|\theta_{1}-\theta_{2}\|.$
Assumption 1 states that the operators $F$, $G$, and $\omega^{\star}$ are Lipschitz continuous and the function $h$ has Lipschitz gradients. This assumption is very common among works on two-time-scale SA (Zeng et al., 2024; Doan, 2022; Hong et al., 2023), and holds true in the RL applications to be discussed in Section 4. Without any loss of generality, we assume $L\geq 1$. We slightly abuse the notation to write
$F(\theta,\omega)\triangleq\mathbb{E}_{X\sim\xi}[F(\theta,\omega,X)],\quad G(\theta,\omega)\triangleq\mathbb{E}_{X\sim\xi}[G(\theta,\omega,X)].$
Assumption 1 obviously implies
$\left\|F(\theta_{1},\omega_{1})-F(\theta_{2},\omega_{2})\right\|\leq L\left(\|\theta_{1}-\theta_{2}\|+\|\omega_{1}-\omega_{2}\|\right),$
$\left\|G(\theta_{1},\omega_{1})-G(\theta_{2},\omega_{2})\right\|\leq L\left(\|\theta_{1}-\theta_{2}\|+\|\omega_{1}-\omega_{2}\|\right).$
Assumption 2 (Bounded Variance)
There exists a positive constant $B$ such that for all $\theta\in\mathbb{R}^{d}$, $\omega\in\mathbb{R}^{r}$, and $X\in{\cal X}$
$\mathbb{E}_{X\sim\xi}[\|F(\theta,\omega,X)-F(\theta,\omega)\|^{2}]\leq B,\quad\mathbb{E}_{X\sim\xi}[\|G(\theta,\omega,X)-G(\theta,\omega)\|^{2}]\leq B.$
The second assumption states that the operators $F$ and $G$ have bounded variance. This assumption is again standard in the literature on stochastic gradient algorithms under i.i.d. noise (Rakhlin et al., 2012; Shamir and Zhang, 2013; Hazan and Kale, 2014). We use ${\cal H}_{k}=\{X_{0},X_{1},\cdots,X_{k}\}$ to denote the randomness information observed until iteration $k$.
Assumption 3 (Strong Monotonicity)
There exists a constant $\mu_{G}>0$ such that
$\left\langle\mathbb{E}_{X\sim\xi}[G(\theta,\omega,X)],\omega-\omega^{\star}(\theta)\right\rangle\geq\mu_{G}\|\omega-\omega^{\star}(\theta)\|^{2},\quad\forall\theta\in\mathbb{R}^{d},\omega\in\mathbb{R}^{r}.$ (12)
Assumption 3 requires the operator $G$ to be strongly monotone in expectation. If $G$ is the gradient of a function $\zeta$, i.e. there exists a function $\zeta:\mathbb{R}^{d}\times\mathbb{R}^{r}\rightarrow\mathbb{R}$ such that $\nabla_{\omega}\zeta(\theta,\omega)=G(\theta,\omega)$ for all $\theta,\omega$, then (12) amounts to the strong convexity of $\zeta$. Again, this assumption can be verified to hold in all RL applications discussed in Section 4.
To simplify notations we introduce the residual variables, which will all be driven to 0
$\begin{gathered}\Delta f_{k}=f_{k}-F\left(\theta_{k},\omega_{k}\right),\,\,\Delta g_{k}=g_{k}-G\left(\theta_{k},\omega_{k}\right),\,\,y_{k}=\|\omega_{k}-\omega^{\star}(\theta_{k})\|^{2}.\end{gathered}$ (14)
3.1 Strong Convexity
We first analyze the convergence of Algorithm 1 when the upper level objective is strongly convex, under the following step sizes
$\lambda_{k}=\frac{c_{\lambda}}{k+\tau+1},\quad\alpha_{k}=\frac{c_{\alpha}}{k+\tau+1},\quad\beta_{k}=\frac{c_{\beta}}{k+\tau+1}.$ (15)
Here $c_{\alpha}$, $c_{\beta}$, $c_{\alpha}$, and $\tau\geq 1$ are properly selected constants depending only on the structure of $h$ and $G$. Their exact choices are presented in Appendix A.
Assumption 4 (Strong Convexity)
The upper level objective $h$ is strongly convex, i.e. there exists a constant $\mu_{h}>0$ such that
$\left\langle\nabla h(\theta_{1})-\nabla h(\theta_{2}),\theta_{1}-\theta_{2}\right\rangle\geq\mu_{h}\|\theta_{1}-\theta_{2}\|^{2},\quad\forall\theta_{1},\theta_{2}\in\mathbb{R}^{d}.$ (16)
Theorem 1
We recall the definition of residual variables in (14) and denote $z_{k}=\|\theta_{k}-\theta^{\star}\|^{2}$. Under Assumptions 1-4 and the step sizes in (15), the iterates of Algorithm 1 satisfy for all $k$
$\mathbb{E}[z_{k+1}]\leq\frac{B\tau^{2}}{k+\tau+1}+\frac{\tau^{2}}{(k+\tau+1)^{2}}\left(\|\Delta f_{0}\|^{2}+\|\Delta g_{0}\|^{2}+z_{0}+y_{0}\right).$ (17)
The theorem states that when $h$ is strongly convex the iterate of Algorithm 1 converge to the globally optimal solution $\theta^{\star}$ in the sense of $\|\theta_{k}-\theta^{\star}\|^{2}$ with a rate of ${\cal O}(1/k)$, which matches the complexity of the standard SGD under the oracle $H$ in (8) (Rakhlin et al., 2012). This rate is much better than the best-known complexity of the standard two-time-scale SA algorithm established in Hong et al. (2023); Zeng et al. (2024), which is ${\cal O}(1/k^{2/3})$. (1)-(3) under a strongly convex $h$ abstracts the objective of gradient-based temporal difference learning in RL, which we present in Section 4.1.
Remark 2
We note that Markovian/i.i.d. sampling results in a difference in the analysis of $f_{k},g_{k}$ in Lemma 2 and 3. In the proof of Lemma 2 in Appendix A.3 for example, it affects how we handle the cross term
$b_{k}=\langle\Delta f_{k},F(\theta,\omega_{k},X_{k})-F(\theta,\omega_{k})\rangle\Rightarrow\left\{\begin{array}[]{ll}\mathbb{E}[b_{k}]=0&\text{under i.i.d samples}\\ \mathbb{E}[b_{k}]\neq 0&\text{under Markovian samples}\end{array}\right.$
When $X_{k}$ is from a Markov chain with $\xi$ as the stationary distribution, $\mathbb{E}[b_{k}]$ captures the gap between $\xi$ and the distribution of $X_{k}$. This gap decays exponentially, under the assumption that the Markov chain is geometrically ergodic, which is a standard assumption in the literature (Xiong et al., 2021; Chen et al., 2022b; Zeng et al., 2022). Using techniques first developed in Srikant and Ying (2019), we can show that $\mathbb{E}[\|b_{k}\|]\leq{\cal O}(\alpha_{k}\log(1/\alpha_{k}))$. The rest of the proof of Theorem 1 can proceed without modification, and eventually we have $\mathbb{E}[z_{k+1}]\leq{\cal O}(\log(k+1)/(k+1))$, only differing from (17) by a logarithm factor. The same extension can be made similarly on the results to be presented later in this section for other types of functions.
3.2 Non-Convexity with Polyak-Lojasiewicz Condition
We now shift to the non-convex regime. The optimal solution to (1) may not be unique without the strong convexity assumption. Let $\Theta^{\star}\subseteq\mathbb{R}^{d}$ denote the set of optimal solutions to (1) and $h^{\star}$ denote the optimal function value. We study the complexity of Algorithm 1 when the upper-level objective function of (1) satisfies the PL condition below.
Assumption 5 (Polyak-Lojasiewicz Condition)
There exists a constant $\mu_{h}>0$ such that
$\frac{1}{2}\|\nabla h(\theta)\|^{2}\geq\mu_{h}\left(h(\theta)-h^{\star}\right),\quad\forall\theta\in\mathbb{R}^{d}.$
We study Algorithm 1 under the following choice of step sizes
$\lambda_{k}=\frac{c_{\lambda}}{k+\tau+1},\quad\alpha_{k}=\frac{c_{\alpha}}{k+\tau+1},\quad\beta_{k}=\frac{c_{\beta}}{k+\tau+1},$
where $c_{\alpha}$, $c_{\beta}$, $c_{\lambda}$, and $\tau\geq 1$ again need to be properly selected according to the structure of $h$ and $G$. Detailed choice of the parameters can be found in Appendix B.
Theorem 2
We recall the residual variables defined in (14) and denote $x_{k}=h(\theta_{k})-h^{\star}$. Under Assumptions 1-3, Assumption 5, and the step sizes above, the iterates of Algorithm 1 satisfy for all $k$
$\mathbb{E}[x_{k+1}]$ $\leq\frac{B\tau^{2}}{k+\tau+1}+\frac{\tau^{2}}{(k+\tau+1)^{2}}\left(\|\Delta f_{0}\|^{2}+\|\Delta g_{0}\|^{2}+x_{0}+y_{0}\right).$
The theorem states that under the PL condition $h(\theta_{k})$ converges with rate ${\cal O}(1/k)$ to the globally optimal function value $h^{\star}$. This again outperforms the standard two-time-scale SA algorithm which has a complexity of ${\cal O}(1/k^{2/3})$. Functions observing the PL condition include the policy optimization objectives in the linear-quadratic regulator and entropy-regularized MDP, which are the subjects of discussion in Section 4.2 and 4.3.
3.3 General Nonconvexity
For non-convex functions without additional structure, we consider step sizes that decay as $1/\sqrt{k}$
$\lambda_{k}=\frac{1}{4(k+1)^{1/2}},\quad\alpha_{k}=\frac{\alpha_{0}}{(k+1)^{1/2}},\quad\beta_{k}=\frac{\beta_{0}}{(k+1)^{1/2}},$ (18)
with $\alpha_{0},\beta_{0}$ chosen such that $\alpha_{0}\leq\beta_{0}$, $\beta_{0}\leq\frac{1}{72L}$, $\beta_{0}\leq\lambda_{0}$, $\alpha_{0}\leq\frac{1}{72L^{2}}$.
Theorem 3
Under Assumptions 1-3 and the step sizes in (18), we have for all $k$
$\min_{t\leq k}\mathbb{E}\left[\|\nabla h(\theta_{t})\|^{2}\right]$ $\leq\frac{32}{\alpha_{0}(k+1)^{1/2}}\left(\|\Delta f_{0}\|^{2}+\|\Delta g_{0}\|^{2}+(h(\theta_{0})-h^{\star})+y_{0}\right)+\frac{4\log(k+2)}{\alpha_{0}(k+1)^{1/2}}.$
The convergence in the non-convex case is not global, but rather to a stationary point of $h$. The state-of-the-art convergence rate of the standard two-time-scale SA algorithm is ${\cal O}(1/k^{2/5})$ (Hong et al., 2023; Zeng et al., 2024), which we improve to ${\cal O}(1/k^{1/2})$. Applications of the optimization framework for non-convex functions include gradient-based policy evaluation under non-linear function approximation which we present in Section 4.1 and sample-based policy optimization under the standard MDP (Wu et al., 2020; Olshevsky and Gharesifard, 2023; Chen and Zhao, 2024).
4 Motivating Applications
4.1 Gradient-Based Policy Evaluation in Discounted-Reward MDP
Temporal difference learning with gradient correction (TDC) is a popular algorithm for policy evaluation in the setting where samples are collected by a behavior policy different from the one to be evaluated. Originally designed to work with linear function approximation (Sutton et al., 2008, 2009), this method has later been extended to the non-linear function approximation setting in Maei et al. (2009).
Consider a Markov decision process (MDP) characterized by ${\cal M}=({\cal S},{\cal A},{\cal P},r,\gamma)$, where ${\cal S}$, ${\cal A}$ are the state and action space, ${\cal P}:{\cal S}\times{\cal A}\rightarrow\Delta_{{\cal S}}$ is the transition kernel, $r:{\cal S}\times{\cal A}\rightarrow[0,1]$ is the reward function, and $\gamma\in(0,1)$ is the discount factor. In the linear function approximation case, each state $s\in{\cal S}$ is associated with a feature vector $\phi(s)\in\mathbb{R}^{d}$. We denote by $\Phi$ the stacked feature matrix
$\Phi=\left[\begin{array}[]{ccc}-&\phi(s_{1})^{T}&-\\ \cdots&\cdots&\cdots\\ -&\phi(s_{|{\cal S}|})^{T}&-\end{array}\right]\in\mathbb{R}^{|{\cal S}|\times d}.$
Policy evaluation studies finding the value function of a policy $\pi\in\Delta_{{\cal A}}^{{\cal S}}$. Given a parameter $\theta\in\mathbb{R}^{d}$, the approximated value function is $\Phi\theta\in\mathbb{R}^{|{\cal S}|}$. Let $\Pi$ denote the orthogonal projection onto the subspace spanned by $\Phi$. The objective of TDC is to find a parameter $\theta$ that minimizes the mean square projected Bellman error
$\min_{\theta}J(\theta)\triangleq\|\Phi\theta-\Pi T^{\pi}\Phi\theta\|_{\mu_{\pi}}^{2},$ (19)
where $\mu_{\pi}$ denotes the stationary distribution of states under policy $\pi$ and $\|v\|_{\mu_{\pi}}^{2}=\sum_{s\in{\cal S}}\mu_{\pi}(s)v(s)^{2}$ for any $v\in\mathbb{R}^{|{\cal S}|}$.
As noted in Sutton et al. (2009), unbiased stochastic gradients of $J$ cannot be directly obtained using samples. Instead, we need to introduce an auxiliary variable $\omega\in\mathbb{R}^{d}$ and solve
$\nabla_{\theta}J(\theta)=-2\mathbb{E}_{\pi}[r(s,a)+\gamma\phi(s^{\prime})^{\top}\theta\phi(s)-\phi(s)^{\top}\theta\phi(s)]+2\gamma\mathbb{E}_{\pi}[\phi(s^{\prime})\phi(s)^{\top}]\,\omega=0,$ (20)
$\mathbb{E}_{\pi}[\phi(s)\phi(s)^{\top}]\,\,\omega=\mathbb{E}_{\pi}[r(s,a)+\gamma\phi(s^{\prime})^{\top}\theta\phi(s)-\phi(s)^{\top}\theta\phi(s)].$ (21)
It is straightforward to see that the objective (20)-(21) and stochastic oracles of the TDC algorithm exactly matches (1)-(3), where the operator in the lower level problem (21) is strongly monotone and the upper level objective is strongly convex. As a result, Algorithm 1 specializes to a variant of the TDC algorithm guaranteed to converge with rate $\widetilde{{\cal O}}(1/k)$, which matches the best known convergence rate of TDC under linear function approximation (Haque et al., 2023). When non-linear function approximation is considered, TDC solves a system of equations that resemble (20)-(21), where the lower level operator is still strongly monotone but the upper level objective is non-convex (Maei et al., 2009). The best known convergence rate for TDC under non-linear function approximation is $\widetilde{{\cal O}}(1/k^{-1/2})$ to a stationary point of $J$ (Wang et al., 2021), which we again match by our result in Theorem 3.
4.2 Policy Optimization for Linear-Quadratic Regulators
Linear-quadratic regulator (LQR) studies finding an optimal control sequence $\{u_{k}\}$ that stabilizes a linear system with the minimum cost
$\{u_{k}^{\star}\}\,\,=\,\,\min_{\{u_{k}\}}\,&\lim_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}\Big{[}\sum_{k=0}^{T}\left(x_{k}^{\top}Qx_{k}+u_{k}^{\top}Ru_{k}\right)\mid x_{0}\Big{]}\\ \text{subject to}\quad&x_{k+1}=Ax_{k}+Bu_{k}+w_{k},\end{aligned}$ (22)
where $x_{k}\in\mathbb{R}^{d_{1}}$, $u_{k}\in\mathbb{R}^{d_{2}}$ are the state and the control variables, $w_{k}\in\mathbb{R}^{d_{1}}$ is the system noise, $A\in\mathbb{R}^{d_{1}\times d_{1}}$ and $B\in\mathbb{R}^{d_{1}\times d_{2}}$ are the transition matrices, and $Q\in\mathbb{R}^{d_{1}\times d_{1}},R\in\mathbb{R}^{d_{2}\times d_{2}}$ are positive-definite cost matrices. We assume that $\{w_{k}\}$ is independently and identically distributed according to $N(0,\Psi)$ for some covariance matrix $\Psi\in\mathbb{R}^{d_{1}\times d_{1}}$.
It is a classic result in the control literature that the optimal control sequence $\{u_{k}^{\star}\}$ is a time-invariant linear function of the state
$u_{k}^{\star}=-K^{\star}x_{k},$ (23)
for a matrix $K^{\star}\in\mathbb{R}^{d_{2}\times d_{1}}$ as a function of $A,B,Q,R$ (Bertsekas, 2012). Exploiting this fact allows us to re-formulate the LQR objective as finding the optimal control matrix $K$, which results in an optimization problem of the following form
$\min_{K}\quad$ $\operatorname{trace}(P_{K}\Psi))$
subject to $P_{K}=Q+K^{\top}RK+(A-BK)^{\top}P_{K}(A-BK).$ (24)
To encourage exploration, we consider the regularized objective with $\sigma\geq 0$ and $\Psi_{\sigma}=\Psi+\sigma^{2}BB^{\top}$
[…]
which is known to have the same optimal solution as the unregularized problem (Gravell et al., 2020).
We are interested in finding the solution to (25) under unknown transition matrices $A$ and $B$, by taking a gradient-based approach. The natural gradient of $J$ with respect to $K$, which we denote by ~$\widetilde{\nabla}_{K}J$ and which can be regarded as $\nabla_{K}J$ on a slightly transformed system of coordinates, has the following closed-form expression
~$\widetilde{\nabla}_{K}J=2\left(R+B^{\top}P_{K}B\right)K-2B^{\top}P_{K}A.$
To obtain a reliable stochastic estimate of ~$\widetilde{\nabla}_{K}J$ the key is to estimate $R+B^{\top}P_{K}B$ and $B^{\top}P_{K}A$. For simplicity of notation, we define
$\Omega_{K}=\left(\begin{array}[]{cc}\Omega_{K}^{11}&\Omega_{K}^{12}\\ \Omega_{K}^{21}&\Omega_{K}^{22}\end{array}\right)=\left(\begin{array}[]{cc}Q+A^{\top}P_{K}A&A^{\top}P_{K}B\\ B^{\top}P_{K}A&R+B^{\top}P_{K}B\end{array}\right),$ (30)
of which $R+B^{\top}P_{K}B$ and $B^{\top}P_{K}A$ are sub-matrices. We also define $\text{svec}(\cdot)$ as the operation to vectorize the upper triangular sub-matrix of a given symmetric matrix with the off-diagonal entries weighted by $\sqrt{2}$. We further define $\phi(x,u)=\text{svec}(\left[x^{\top},u^{\top}\right]^{\top}\left[x^{\top},u^{\top}\right])$ for any $x\in\mathbb{R}^{d_{1}},u\in\mathbb{R}^{d_{2}}$, which can be seen as the feature vector for the system in state $(x,u)$. It is shown in Yang et al. (2019) that $\Omega_{K}$ and $J(K)$ jointly satisfy
$\mathbb{E}_{x\sim\mu_{K},u\sim N(-Kx,\sigma^{2}I)}\left[M_{x,u,x^{\prime},u^{\prime}}\right]\left[\begin{array}[]{c}J(K)\\ \text{svec}(\Omega_{K})\end{array}\right]=\mathbb{E}_{x\sim\mu_{K},u\sim N(-Kx,\sigma^{2}I)}\left[c_{x,u}\right],$ (33)
where the matrix $M_{x,u,x^{\prime},u^{\prime}}$ and vector $c_{x,u}$ are
$M_{x,u,x^{\prime},u^{\prime}}=\left[\begin{array}[]{cc}1&0\\ \phi(x,u)&\phi(x,u)\left[\phi(x,u)-\phi\left(x^{\prime},u^{\prime}\right)\right]^{\top}\end{array}\right],\,\,c_{x,u}=\left[\begin{array}[]{c}x^{\top}Qx+u^{\top}Ru\\ \left(x^{\top}Qx+u^{\top}Ru\right)\phi(x,u)\end{array}\right].$
Equation (33) maps to the lower level problem (3). Under the assumption that $K$ is a stable controller with respect to $A$ and $B$ (Yang et al., 2019), the linear operator $\mathbb{E}[M_{x,u,x^{\prime},u^{\prime}}]$ is strongly monotone (Yang et al., 2019). The upper level problem that corresponds to (2) is to solve ~$\widetilde{\nabla}_{K}J=0$, which is known to observe the PL condition (Fazel et al., 2018). This means that we can apply Algorithm 1 to this problem, which specializes to an single-looped “actor-critic” algorithm with a finite-time convergence rate of $\widetilde{{\cal O}}(1/k)$ to the globally optimal solution measured by $J(K_{k})-J(\theta^{\star})$. The complexity of the current state-of-the-art single-looped algorithm for LQR is $\widetilde{{\cal O}}(1/k^{2/3})$, which we improve over. We also note the recent few works on nested-loop actor-critic algorithms that rely on batched multi-trajectory samples (Ju et al., 2022; Zhou and Lu, 2023). These algorithms achieve the complexity $\widetilde{{\cal O}}(1/k)$ as well, but in practice are much less convenient to implement than our algorithm which is single-looped and only requires a single online trajectory of samples.
Algorithm 2 Fast Single-Loop Actor-Critic Algorithm for LQR
1:  Initialization: Control matrix $K_{0}$, auxiliary variable $\hat{\Omega}_{0},\hat{J}_{0}$, gradient estimation variable $f_{0},g_{0}^{J},g_{0}^{\Omega}$, step size sequences $\{\alpha_{k},\beta_{k},\lambda_{k}\}$
2:  Sample an initial state $x_{0}\sim{\cal D}$, takes an initial control $u_{0}\sim N(-K_{0}x_{0},\sigma^{2}I)$, and observe cost $c_{0}=x_{0}^{\top}Qx_{0}+u_{0}^{\top}Ru_{0}$ and next state $x_{1}\sim N(Ax_{0}+Bu_{0},\Psi)$
3:  for $k=0,1,2,...$ do
4:     Take the control $u_{k+1}\sim N(-K_{k}x_{k+1},\sigma^{2}I)$. Observe cost $c_{k+1}=x_{k+1}^{\top}Qx_{k+1}+u_{k+1}^{\top}Ru_{k+1}$ and next state $x_{k+2}\sim N(Ax_{k+1}+Bu_{k+1},\Psi)$
5:     Control matrix update:
$K_{k+1}=K_{k}-\alpha_{k}f_{k}$
6:     Auxiliary variable update:
$\hat{J}_{k+1}=\hat{J}_{k}-\beta_{k}g_{k}^{J},\quad\hat{\Omega}_{k+1}=\hat{\Omega}_{k}-\beta_{k}g_{k}^{\Omega}$
7:     Gradient estimation variable update:
$f_{k+1}$ $=(1-\lambda_{k})f_{k}+\lambda_{k}(\hat{\Omega}_{k}^{22}K_{k}-\hat{\Omega}_{k}^{21})$
$g_{k+1}^{J}$ $=(1-\lambda_{k})g_{k}^{J}+\lambda_{k}(\hat{J}_{k}-x_{k}^{\top}Qx_{k}-u_{k}^{\top}Ru_{k})$
$g_{k+1}^{\Omega}$ $=(1-\lambda_{k})g_{k}^{\Omega}-\lambda_{k}\left[\begin{array}[]{c}x_{k}\\ u_{k}\end{array}\right]\left[\begin{array}[]{c}x_{k}\\ u_{k}\end{array}\right]^{\top}\left(\left[x_{k}^{\top}\,\,\,u_{k}^{\top}\right]\hat{\Omega}_{k}\left[\begin{array}[]{c}x_{k}\\ u_{k}\end{array}\right]+\hat{J}_{k}-x_{k}^{\top}Qx_{k}-u_{k}^{\top}Ru_{k}\right)$
8:  end for
4.3 Policy Optimization for Entropy-Regularized MDPs
The optimization framework (1)-(3) also applies to the policy optimization problem in the entropy-regularized MDP. Consider the MDP ${\cal M}=({\cal S},{\cal A},{\cal P},r,\gamma)$ introduced in Section 4.1. The regularized value function of a policy $\pi\in\Delta_{{\cal A}}^{{\cal S}}$ is
$V_{\tau}^{\pi}(s)$ $=\mathbb{E}_{a_{k}\sim\pi(\cdot\mid s_{k}),s_{k+1}\sim{\cal P}(\cdot\mid s_{k},a_{k})}\Big{[}\sum_{k=0}^{\infty}\gamma^{k}\big{(}r(s_{k},a_{k})-\tau\log\pi(a_{k}\mid s_{k})\big{)}\mid s_{0}=s\Big{]},$
where $\tau$ is a positive regularization weight. Under initial state distribution $\rho\in\Delta_{{\cal S}}$, the expected cumulative reward under policy $\pi$ is $J_{\tau}(\pi)\triangleq\mathbb{E}_{s\sim\rho}[V_{\tau}^{\pi}(s)]$. We consider the tabular setting where the policy is represented through the softmax function by a parameter $\theta\in\mathbb{R}^{|{\cal S}|\times|{\cal A}|}$
$\pi_{\theta}(a\mid s)$ $=\frac{\exp\left(\theta_{s,a}\right)}{\sum_{a^{\prime}\in{\cal A}}\exp\left(\theta_{s,a^{\prime}}\right)}.$
Policy optimization studies finding the parameter $\theta^{\star}$ that maximizes the cumulative reward $J_{\tau}(\pi_{\theta})$. By the first-order condition, this is equivalent to solving
$\nabla_{\theta}J_{\tau}(\pi_{\theta})=\frac{1}{1-\gamma}\mathbb{E}_{\pi_{\theta}}\big{[}\left(r(s,a)-\tau\log\pi_{\theta}(a\mid s)+\gamma V_{\tau}^{\pi_{\theta}}(s^{\prime})-V_{\tau}^{\pi_{\theta}}(s)\right)\nabla_{\theta}\log\pi_{\theta}(a\mid s)\big{]}=0.$
Evaluating the gradient requires computing the regularized value function $V_{\tau}^{\pi_{\theta}}\in\mathbb{R}^{|{\cal S}|}$, which satisfies the regularized Bellman equation
$\mathbb{E}_{\pi_{\theta}}\Big{[}r(s,a)-\tau\log\pi_{\theta}(a\mid s)+\gamma V_{\tau}^{\pi_{\theta}}(s^{\prime})-V_{\tau}^{\pi_{\theta}}(s)\Big{]}=0.$ (34)
This objective again falls under the general two-time-scale optimization framework, with the operator in lower level problem (34) being strongly monotone. The upper level objective function $J_{\tau}$ satisfies the PL condition (Mei et al., 2020). Applying Algorithm 1 to this problem results in a fast actor-critic algorithm (Algorithm 2) that is able to reduce the optimality gap $J_{\tau}(\theta_{k})-J_{\tau}(\theta^{\star})$ with rate $\widetilde{{\cal O}}(1/k)$. The analysis in Zeng et al. (2024) implies that the standard two-time-scale algorithm (7) solves this problem with a complexity of $\widetilde{{\cal O}}(1/k^{2/3})$, which we vastly improve over.
5 Numerical Simulations
We experimentally verify that Algorithm 1 converges faster than (7) in solving 1) a policy evaluation problem under linear function approximation, and 2) policy optimization for the LQR.
Policy evaluation problem under linear function approximation. As we have discussed in Section 4.1, TDC is the state-of-the-art policy evaluation algorithm under off-policy samples which enjoys a convergence rate of $\widetilde{{\cal O}}(1/k)$ under linear function approximation. Numerically we evaluate Algorithm 1 against TDC on the task of evaluating the value function of a randomly generated policy in a randomly generated environment with $|{\cal S}|=|{\cal A}|=50$ and the feature dimension $d=10$. Figure 2 shows the empirical performance gap between the standard TDC and Algorithm 1, though both algorithm have the same complexity in theory. For comparison we also include TD(0) learning under linear function approximation. The simulation results in Figure 2 numerically confirms the superiority of the proposed algorithm. Details of the experimental setup can be found in Appendix D.
Figure 1: Fast TDC Algorithm for Random Policy Evaluation
Figure 2: Performance of Fast Actor-Critic Algorithm for LQR
Policy optimization for LQR. We experiment with the small-dimensional environment considered in Zeng et al. (2024) where the transition and costs matrices are chosen to be
$A=\left[\begin{array}[]{ccc}0.5&0.01&0\\ 0.01&0.5&0.01\\ 0&0.01&0.5\end{array}\right],\quad B=\left[\begin{array}[]{ll}1&0.1\\ 0&0.1\\ 0&0.1\end{array}\right],\quad Q=I_{3},\quad R=I_{2}.$ (41)
We apply Algorithm 2 to this problem along with online actor-critic algorithm studied in Zeng et al. (2024), which is equivalent to (7) applied to this context. In Figure 2, we plot the error decay under the two algorithms, which again shows that the proposed algorithm enjoys faster convergence.
Disclaimer
This paper was prepared for informational purposes in part by the Artificial Intelligence Research group of JP Morgan Chase & Co and its affiliates (“JP Morgan”), and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
Acknowledgement
This work was supported by the National Science Foundation under awards ECCS-CAREER-2339509 and CCF-2343599.
Appendix A Analysis for Strongly Convex Functions
The exact choice of the step size parameters is made such that $\alpha_{k}\leq\beta_{k}\leq\lambda_{k}\leq 1/4$, $c_{\alpha}\geq\frac{8}{\mu_{h}}$, and
$\alpha_{k}\leq\min\{\frac{\lambda_{k}}{\mu_{h}},\frac{\mu_{h}}{6(L+1)^{4}},\frac{8\mu_{G}}{\mu_{h}}\beta_{k},\frac{\mu_{h}\mu_{G}}{152(L+1)^{6}}\beta_{k},\frac{\mu_{G}\beta_{k}}{8(8L^{4}+\frac{9L^{2}}{\mu_{G}})},\frac{\lambda_{k}}{4}(48L^{2}+\frac{10L}{\mu_{G}})^{-1}\},$
$\beta_{k}\leq\min\{\frac{1}{240(L+1)^{6}},\frac{1}{\mu_{G}},\frac{\lambda_{k}}{4}(56L^{2}+\frac{9}{2\mu_{h}})^{-1},\frac{\lambda_{k}}{4}(48L^{2}+\frac{10L}{\mu_{G}})^{-1},\frac{\mu_{G}}{168L^{4}}\},$
which requires $\beta_{0}$ to be selected sufficiently small and $\tau$ sufficiently large. Note that parameters $c_{\alpha}$, $\beta_{0}$, and $\tau$ always exist and only depend on the structural constants of $h$ and $G$.
The proof of Theorem 1 is built on the lemmas below. Their proofs can be found at the end of the section.
Lemma 1
Under Assumption 1-3, the sequence of variables $\{\theta_{k},\omega_{k},f_{k},g_{k}\}$ satisfy for all $k$
$\|f_{k}\|$ $\leq\|\Delta f_{k}\|+L\sqrt{y_{k}}+L(L+1)\sqrt{z_{k}},$
$\|g_{k}\|$ $\leq\|\Delta g_{k}\|+L\sqrt{y_{k}},$
$\|\theta_{k+1}-\theta_{k}\|$ $\leq\alpha_{k}\left(\|\Delta f_{k}\|+L\sqrt{y_{k}}+L(L+1)\sqrt{z_{k}}\right),$
$\|\omega_{k+1}-\omega_{k}\|$ $\leq\beta_{k}\left(\|\Delta g_{k}\|+L\sqrt{y_{k}}\right).$
Lemma 2
Suppose that the step size $\lambda_{k}$ satisfy $\lambda_{k}\leq 1/4$ for all $k$. Then, under Assumption 1-3, we have the following bound on the iteration-wise convergence of $\mathbb{E}[\|\Delta f_{k}\|^{2}]$.
$\mathbb{E}[\|\Delta f_{k+1}\|^{2}]$ $\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6(L+1)^{6}\alpha_{k}^{2}z_{k}\right]+2B\lambda_{k}^{2}.$
Lemma 3
Suppose that the step size $\lambda_{k}$ satisfy $\lambda_{k}\leq 1/4$ for all $k$. Then, under Assumption 1-3, we have the following bound on the iteration-wise convergence of $\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\mathbb{E}[\|\Delta g_{k+1}\|^{2}]$ $\leq(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6(L+1)^{6}\alpha_{k}^{2}z_{k}\right]+2B\lambda_{k}^{2}.$
Lemma 4
Suppose the step size $\alpha_{k}$ is chosen such that $\alpha_{k}\leq\frac{\mu_{h}}{6(L+1)^{4}}$ and $\alpha_{k}\leq\frac{1}{2\mu_{h}}$ for all $k$. Under Assumption 1-4, we have
$z_{k+1}\leq(1-\frac{\mu_{h}\alpha_{k}}{4})z_{k}+\frac{9L^{2}\alpha_{k}}{\mu_{h}}y_{k}+\frac{9\alpha_{k}}{2\mu_{h}}\|\Delta f_{k}\|^{2}.$
Lemma 5
Suppose that the step sizes satisfy $\alpha_{k}\leq\beta_{k}\leq\frac{1}{72(L+1)^{6}}$ and $\beta_{k}\leq 1/\mu_{G}$ for all $k$. Then, under Assumption 1-3, we have
$y_{k+1}$ $\leq(1-\mu_{G}\beta_{k})y_{k}-\left(\frac{\mu_{G}}{4}\beta_{k}-8L^{4}\alpha_{k}-L^{2}\beta_{k}^{2}\right)y_{k}$
$\hskip 20.0pt+\left(\frac{19(L+1)^{6}\alpha_{k}^{2}}{\mu_{G}\beta_{k}}+3(L+1)^{6}\alpha_{k}\beta_{k}\right)z_{k}+\left(\frac{8\beta_{k}}{\mu_{G}}+2L\beta_{k}^{2}\right)\|\Delta g_{k}\|^{2}+8L\alpha_{k}\|\Delta f_{k}\|^{2}.$
A.1 Proof of Theorem 1
We consider the Lyapunov function
$V_{k}\triangleq\mathbb{E}[\|\Delta f_{k}\|^{2}+\|\Delta g_{k}\|^{2}+z_{k}+y_{k}].$
Combining Lemma 2-5, we get
$V_{k+1}$ $=\mathbb{E}[\|\Delta f_{k+1}\|^{2}+\|\Delta g_{k+1}\|^{2}+z_{k+1}+y_{k+1}]$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6(L+1)^{6}\alpha_{k}^{2}z_{k}\right]+2B\lambda_{k}^{2}$
$\hskip 20.0pt+(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6(L+1)^{6}\alpha_{k}^{2}z_{k}\right]+2B\lambda_{k}^{2}$
$\hskip 20.0pt+\mathbb{E}[(1-\frac{\mu_{h}\alpha_{k}}{4})z_{k}+\frac{9L^{2}\alpha_{k}}{\mu_{h}}y_{k}+\frac{9\alpha_{k}}{2\mu_{h}}\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+\mathbb{E}[(1-\mu_{G}\beta_{k})y_{k}-\left(\frac{\mu_{G}}{4}\beta_{k}-8L^{4}\alpha_{k}-L^{2}\beta_{k}^{2}\right)y_{k}$
$\hskip 20.0pt+\left(\frac{19(L+1)^{6}\alpha_{k}^{2}}{\mu_{G}\beta_{k}}+3(L+1)^{6}\alpha_{k}\beta_{k}\right)z_{k}+\left(\frac{8\beta_{k}}{\mu_{G}}+2L\beta_{k}^{2}\right)\|\Delta g_{k}\|^{2}+8L\alpha_{k}\|\Delta f_{k}\|^{2}]$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{48L^{2}\beta_{k}^{2}}{\lambda_{k}}-\frac{9\alpha_{k}}{2\mu_{h}}-8L\alpha_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{48L^{2}\beta_{k}^{2}}{\lambda_{k}}-\frac{8\beta_{k}}{\mu_{G}}-2L\beta_{k}^{2})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+(1-\frac{\mu_{h}\alpha_{k}}{8})\mathbb{E}[z_{k}]-(\frac{\mu_{h}\alpha_{k}}{8}-15(L+1)^{6}\alpha_{k}\beta_{k}-\frac{19(L+1)^{6}\alpha_{k}^{2}}{\mu_{G}\beta_{k}})\mathbb{E}[z_{k}]$
$\hskip 20.0pt+(1-\mu_{G}\beta_{k})\mathbb{E}[y_{k}]-\left(\frac{\mu_{G}}{4}\beta_{k}-8L^{4}\alpha_{k}-21L^{4}\beta_{k}^{2}-\frac{9L^{2}\alpha_{k}}{\mu_{h}}\right)\mathbb{E}[y_{k}]+4B\lambda_{k}^{2}$
$\leq(1-\frac{\mu_{h}\alpha_{k}}{8})V_{k}+4B\lambda_{k}^{2}-(\frac{\lambda_{k}}{4}-\frac{48L^{2}\beta_{k}^{2}}{\lambda_{k}}-\frac{9\alpha_{k}}{2\mu_{h}}-8L\alpha_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt-(\frac{\lambda_{k}}{4}-\frac{48L^{2}\beta_{k}^{2}}{\lambda_{k}}-\frac{8\beta_{k}}{\mu_{G}}-2L\beta_{k}^{2})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt-(\frac{\mu_{h}\alpha_{k}}{8}-15(L+1)^{6}\alpha_{k}\beta_{k}-\frac{19(L+1)^{6}\alpha_{k}^{2}}{\mu_{G}\beta_{k}})\mathbb{E}[z_{k}]$
$\hskip 20.0pt-\left(\frac{\mu_{G}}{4}\beta_{k}-8L^{4}\alpha_{k}-21L^{4}\beta_{k}^{2}-\frac{9L^{2}\alpha_{k}}{\mu_{h}}\right)\mathbb{E}[y_{k}],$ (42)
where the third inequality simplifies the terms with the conditions $\alpha_{k}\leq\frac{\lambda_{k}}{\mu_{h}}$ and $\alpha_{k}\leq\frac{8\mu_{G}\beta_{k}}{\mu_{h}}$.
The terms on the right hand side of (42) except for the first two are non-positive under the step size conditions $\beta_{k}\leq\frac{1}{\mu_{G}}$, $\beta_{k}\leq\frac{\lambda_{k}}{4}(56L^{2}+\frac{9}{2\mu_{h}})^{-1}$, $\beta_{k}\leq\frac{\lambda_{k}}{4}(48L^{2}+\frac{10L}{\mu_{G}})^{-1}$, $\alpha_{k}\leq\frac{\mu_{h}\mu_{G}}{152(L+1)^{6}}\beta_{k}$, $\beta_{k}\leq\frac{1}{240(L+1)^{6}}$, $\beta_{k}\leq\frac{\mu_{G}}{168L^{4}}$, and $\alpha_{k}\leq\frac{\mu_{G}\beta_{k}}{8(8L^{4}+\frac{9L^{2}}{\mu_{G}})}$. This leads to
$V_{k+1}\leq(1-\frac{\mu_{h}\alpha_{k}}{8})V_{k}+4B\lambda_{k}^{2}\leq(1-\frac{2}{k+\tau+1})V_{k}+4B\lambda_{k}^{2},$
where the second inequality follows from $c_{\alpha}\geq\frac{16}{\mu_{h}}$.
Multiplying by $(k+\tau+1)^{2}$ on both sides, we have
$(k+\tau+1)^{2}V_{k+1}$ $\leq(k+\tau+1)(k+\tau-1)V_{k}+4B(k+\tau+1)^{2}\lambda_{k}^{2}$
$\leq(k+\tau)^{2}V_{k}+\frac{B(k+\tau+1)^{2}}{4(k+1)^{2}}$
$\leq(k+\tau)^{2}V_{k}+(\frac{B}{2}+\frac{B\tau^{2}}{2})$
$\leq\tau^{2}V_{0}+B\tau^{2}(k+1)$
Dividing by $(k+\tau+1)^{2}$ now, we have
$V_{k+1}\leq\frac{\tau^{2}V_{0}}{(k+\tau+1)^{2}}+\frac{B\tau^{2}(k+1)}{(k+\tau+1)^{2}}\leq\frac{\tau^{2}V_{0}}{(k+\tau+1)^{2}}+\frac{B\tau^{2}}{(k+\tau+1)},$
which obviously implies
$\mathbb{E}[\|\theta_{k+1}-\theta^{\star}\|^{2}]\leq\frac{\tau^{2}V_{0}}{(k+\tau+1)^{2}}+\frac{B\tau^{2}}{(k+\tau+1)}.$
$\blacksquare$
A.2 Proof of Lemma 1
By definition,
$F(\theta^{\star},\omega^{\star}(\theta^{\star}))=\nabla h(\theta^{\star})=0,$
which implies
$\|f_{k}\|$ $=\|\Delta f_{k}+F(\theta_{k},\omega_{k})-F(\theta_{k},\omega^{\star}(\theta_{k}))+F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta^{\star},\omega^{\star}(\theta^{\star}))\|$
$\leq\|\Delta f_{k}\|+\|F(\theta_{k},\omega_{k})-F(\theta_{k},\omega^{\star}(\theta_{k}))\|+\|F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta^{\star},\omega^{\star}(\theta^{\star}))\|$
$\leq\|\Delta f_{k}\|+L\|\omega_{k}-\omega^{\star}(\theta_{k})\|+L(L+1)\|\theta_{k}-\theta^{\star}\|$
$\leq\|\Delta f_{k}\|+L\sqrt{y_{k}}+L(L+1)\sqrt{z_{k}},$
where the second inequality follows from the Lipschitz continuity of $F$.
Similarly, since $G(\theta,\omega^{\star}(\theta))=0$ for any $\theta$, we can derive
$\|g_{k}\|$ $=\|\Delta g_{k}+G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|$
$\leq\|\Delta g_{k}\|+\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|$
$\leq\|\Delta g_{k}\|+L\sqrt{y_{k}}.$
The bounds on $\|\theta_{k+1}-\theta_{k}\|$ and $\|\omega_{k+1}-\omega_{k}\|$ easily follow from the two inequalities above and (10).
$\blacksquare$
A.3 Proof of Lemma 2
By the update rule in (11), we have
$\Delta f_{k+1}$ $=f_{k+1}-F(\theta_{k+1},\omega_{k+1})$
$=(1-\lambda_{k})f_{k}+\lambda_{k}F(\theta_{k},\omega_{k},X_{k})-F(\theta_{k+1},\omega_{k+1})$
$=(1-\lambda_{k})f_{k}+\lambda_{k}F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})+\lambda_{k}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right)$
$=(1-\lambda_{k})\Delta f_{k}+F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})+\lambda_{k}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right).$
This leads to
$\|\Delta f_{k+1}\|^{2}$
$=(1-\lambda_{k})^{2}\|\Delta f_{k}\|^{2}+\Big{\|}\left(F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\right)+\lambda_{k}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right)\Big{\|}^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\Delta f_{k}^{\top}\left(F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\right)$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta f_{k}^{\top}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right)$
$\leq(1-\lambda_{k})^{2}\|\Delta f_{k}\|^{2}+2\|F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\|^{2}+2\lambda_{k}^{2}\|F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\|^{2}$
$\hskip 20.0pt+\frac{\lambda_{k}}{2}\|\Delta f_{k}\|^{2}+\frac{2}{\lambda_{k}}\left\|F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\right\|^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta f_{k}^{\top}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right)$
$\leq(1-\lambda_{k})\|\Delta f_{k}\|^{2}+2\lambda_{k}^{2}\|F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\|^{2}$
$\hskip 20.0pt+(\lambda_{k}^{2}-\frac{\lambda_{k}}{2})\|\Delta f_{k}\|^{2}+\frac{4}{\lambda_{k}}\left\|F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\right\|^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta f_{k}^{\top}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right),$ (43)
where the last inequality follows from $\lambda_{k}\leq 1$.
The second term on the right hand side of (43) is bounded in expectation, i.e.
$\mathbb{E}_{X_{k}\sim\xi}[\|F(\theta_{k},\omega_{k},X_{k})-F(\theta_{k},\omega_{k})\|^{2}]\leq B.$
In addition, the last term of (43) is zero in expectation since $\mathbb{E}[\Delta f_{k}^{\top}(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k}))\mid{\cal H}_{k-1}]=\Delta f_{k}^{\top}\mathbb{E}[(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k}))\mid{\cal H}_{k-1}]=0$. As a result, we have
$\mathbb{E}[\|\Delta f_{k+1}\|^{2}]$ $\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{2}-\lambda_{k}^{2})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}[\left\|F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\right\|^{2}]+2B\lambda_{k}^{2}.$ (44)
By Lemma 1 and the Lipschitz condition of $F$,
$\|F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\|^{2}$
$\leq 2L^{2}\|\theta_{k+1}-\theta_{k}\|^{2}+2L^{2}\|\omega_{k+1}-\omega_{k}\|^{2}$
$\leq 6L^{2}\alpha_{k}^{2}(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+(L+1)^{4}z_{k})+4L^{2}\beta_{k}^{2}(\|\Delta g_{k}\|^{2}+L^{2}y_{k})$
$\leq 6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6(L+1)^{6}\alpha_{k}^{2}z_{k}.$
Plugging this bound into (44), we get
$\mathbb{E}[\|\Delta f_{k+1}\|^{2}]$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{2}-\lambda_{k}^{2})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6(L+1)^{6}\alpha_{k}^{2}z_{k}\right]+2B\lambda_{k}^{2}$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6(L+1)^{6}\alpha_{k}^{2}z_{k}\right]+2B\lambda_{k}^{2},$
where the last inequality follows from $\lambda_{k}\leq 1/4$.
$\blacksquare$
A.4 Proof of Lemma 3
By the update rule in (11), we have
$\Delta g_{k+1}$ $=g_{k+1}-G(\theta_{k+1},\omega_{k+1})$
$=(1-\lambda_{k})g_{k}+\lambda_{k}G(\theta_{k},\omega_{k},X_{k})-G(\theta_{k+1},\omega_{k+1})$
$=(1-\lambda_{k})g_{k}+\lambda_{k}G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})+\lambda_{k}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right)$
$=(1-\lambda_{k})\Delta g_{k}+G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})+\lambda_{k}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right).$
This leads to
$\|\Delta g_{k+1}\|^{2}$
$=(1-\lambda_{k})^{2}\|\Delta g_{k}\|^{2}+\Big{\|}\left(G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\right)+\lambda_{k}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right)\Big{\|}^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\Delta g_{k}^{\top}\left(G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\right)$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta g_{k}^{\top}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right)$
$\leq(1-\lambda_{k})^{2}\|\Delta g_{k}\|^{2}+2\|G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\|^{2}+2\lambda_{k}^{2}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\|^{2}$
$\hskip 20.0pt+\frac{\lambda_{k}}{2}\|\Delta g_{k}\|^{2}+\frac{2}{\lambda_{k}}\left\|G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\right\|^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta g_{k}^{\top}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right)$
$\leq(1-\lambda_{k})\|\Delta g_{k}\|^{2}+2\lambda_{k}^{2}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\|^{2}$
$\hskip 20.0pt+(\lambda_{k}^{2}-\frac{\lambda_{k}}{2})\|\Delta g_{k}\|^{2}+\frac{4}{\lambda_{k}}\left\|G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\right\|^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta g_{k}^{\top}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right).$ (45)
The second term on the right hand side of (45) is bounded in expectation, i.e.
$\mathbb{E}_{X_{k}\sim\xi}[\|G(\theta_{k},\omega_{k},X_{k})-G(\theta_{k},\omega_{k})\|^{2}]\leq B.$
In addition, the last term of (43) is zero in expectation since $\mathbb{E}[\Delta g_{k}^{\top}(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k}))\mid{\cal H}_{k-1}]=\Delta g_{k}^{\top}\mathbb{E}[(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k}))\mid{\cal H}_{k-1}]=0$. As a result, we have
$\mathbb{E}[\|\Delta g_{k+1}\|^{2}]$ $\leq(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{2}-\lambda_{k}^{2})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}[\left\|G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\right\|^{2}]+2B\lambda_{k}^{2}.$ (46)
By Lemma 1 and the Lipschitz condition of $G$,
$\|G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\|^{2}$
$\leq 2L^{2}\|\theta_{k+1}-\theta_{k}\|^{2}+2L^{2}\|\omega_{k+1}-\omega_{k}\|^{2}$
$\leq 6L^{2}\alpha_{k}^{2}(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+(L+1)^{4}z_{k})+4L^{2}\beta_{k}^{2}(\|\Delta g_{k}\|^{2}+L^{2}y_{k})$
$\leq 6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6(L+1)^{6}\alpha_{k}^{2}z_{k}.$
Plugging this bound into (46), we get
$\mathbb{E}[\|\Delta g_{k+1}\|^{2}]$ $\leq(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{2}-\lambda_{k}^{2})\mathbb{E}[\|\Delta g_{k}\|^{2}]+2B\lambda_{k}^{2}$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6(L+1)^{6}\alpha_{k}^{2}z_{k}\right]$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6(L+1)^{6}\alpha_{k}^{2}z_{k}\right]+2B\lambda_{k}^{2},$
where the last inequality follows from $\lambda_{k}\leq 1/4$.
$\blacksquare$
A.5 Proof of Lemma 4
We have by the update rule in (10) and the definition of the residual variables in (14)
$z_{k+1}$ $=\|\theta_{k+1}-\theta^{\star}\|^{2}$
$=\|\theta_{k}-\alpha_{k}f_{k}-\theta^{\star}\|^{2}$
$=\|\theta_{k}-\alpha_{k}F(\theta_{k},\omega_{k})-\alpha_{k}\Delta f_{k}-\theta^{\star}\|^{2}$
$=\|\theta_{k}-\theta^{\star}-\alpha_{k}F(\theta_{k},\omega_{k})\|^{2}-2\alpha_{k}\langle\theta_{k}-\theta^{\star}-\alpha_{k}F(\theta_{k},\omega_{k}),\Delta f_{k}\rangle+\alpha_{k}^{2}\|\Delta f_{k}\|^{2}.$ (47)
To bound the first term on the right hand side of (47),
$\|\theta_{k}-\theta^{\star}-\alpha_{k}F(\theta_{k},\omega_{k})\|^{2}$
$=\|\theta_{k}-\theta^{\star}-\alpha_{k}\left(F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta^{\star},\omega^{\star}(\theta^{\star}))\right)+\alpha_{k}\left(F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\right)\|^{2}$
$=z_{k}+\alpha_{k}^{2}\|F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta^{\star},\omega^{\star}(\theta^{\star}))\|^{2}+\alpha_{k}^{2}\|F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\|^{2}$
$\hskip 20.0pt-2\alpha_{k}\langle\theta_{k}-\theta^{\star},F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta^{\star},\omega^{\star}(\theta^{\star}))\rangle$
$\hskip 20.0pt+2\alpha_{k}\langle\theta_{k}-\theta^{\star},F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\rangle$
$\hskip 20.0pt-2\alpha_{k}^{2}\langle F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta^{\star},\omega^{\star}(\theta^{\star}),F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\rangle$
$\leq z_{k}+\alpha_{k}^{2}\|F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta^{\star},\omega^{\star}(\theta^{\star}))\|^{2}+\alpha_{k}^{2}\|F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\|^{2}$
$\hskip 20.0pt-2\mu_{h}\alpha_{k}z_{k}+\mu_{h}\alpha_{k}z_{k}+\frac{\alpha_{k}}{\mu_{h}}\|F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\|^{2}$
$\hskip 20.0pt+\alpha_{k}^{2}z_{k}+\alpha_{k}^{2}\|F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\|^{2}$
$\leq(1-\mu_{h}\alpha_{k}+\alpha_{k}^{2})z_{k}+2L^{2}(L+1)^{2}\alpha_{k}^{2}z_{k}+(2L^{2}\alpha_{k}^{2}+\frac{L^{2}\alpha_{k}}{\mu_{h}})y_{k}$
$\leq(1-\mu_{h}\alpha_{k}+3(L+1)^{4}\alpha_{k}^{2})z_{k}+(2L^{2}\alpha_{k}^{2}+\frac{L^{2}\alpha_{k}}{\mu_{h}})y_{k},$
where the first equality follows from $F(\theta^{\star},\omega^{\star}(\theta^{\star}))=0$, the first inequality plugs in (16) and uses the fact that $2\|\boldsymbol{a}\|\|\boldsymbol{b}\|\leq c\|\boldsymbol{a}\|^{2}+\frac{1}{c}\|\boldsymbol{b}\|^{2}$ for any vector $\boldsymbol{a},\boldsymbol{b}$ and scalar $c>0$, and the second inequality follows from the Lipschitz continuity of $F$ and $\omega^{\star}$. Choosing the step size such that $\alpha_{k}\leq\frac{\mu_{h}}{6(L+1)^{4}}$ and $\alpha_{k}\leq\frac{1}{2\mu_{h}}$ results in
$\|\theta_{k}-\theta^{\star}-\alpha_{k}F(\theta_{k},\omega_{k})\|^{2}\leq(1-\frac{\mu_{h}\alpha_{k}}{2})z_{k}+\frac{2L^{2}\alpha_{k}}{\mu_{h}}y_{k}.$ (48)
The second term on the right hand side of (47) can be treated based on (48)
$-2\alpha_{k}\langle\theta_{k}-\theta^{\star}-\alpha_{k}F(\theta_{k},\omega_{k}),\Delta f_{k}\rangle$
$\leq\frac{\mu_{h}\alpha_{k}}{4}\|\theta_{k}-\theta^{\star}-\alpha_{k}F(\theta_{k},\omega_{k})\|^{2}+\frac{4\alpha_{k}}{\mu_{h}}\|\Delta f_{k}\|^{2}$
$\leq\frac{\mu_{h}\alpha_{k}}{4}(1-\frac{\mu_{h}\alpha_{k}}{2})z_{k}+\frac{\mu_{h}\alpha_{k}}{4}\cdot\frac{2L^{2}\alpha_{k}}{\mu_{h}}y_{k}+\frac{4\alpha_{k}}{\mu_{h}}\|\Delta f_{k}\|^{2}$
$\leq(\frac{\mu_{h}\alpha_{k}}{4}-\frac{\mu_{h}^{2}\alpha_{k}^{2}}{8})z_{k}+\frac{L^{2}\alpha_{k}^{2}}{2}y_{k}+\frac{4\alpha_{k}}{\mu_{h}}\|\Delta f_{k}\|^{2}.$ (49)
Plugging (48) and (49) into (47), we get
$z_{k+1}$ $\leq(1-\frac{\mu_{h}\alpha_{k}}{2})z_{k}+\frac{2L^{2}\alpha_{k}}{\mu_{h}}y_{k}$
$\hskip 20.0pt+(\frac{\mu_{h}\alpha_{k}}{4}-\frac{\mu_{h}^{2}\alpha_{k}^{2}}{8})z_{k}+\frac{L^{2}\alpha_{k}^{2}}{2}y_{k}+\frac{4\alpha_{k}}{\mu_{h}}\|\Delta f_{k}\|^{2}+\alpha_{k}^{2}\|\Delta f_{k}\|^{2}$
$\leq(1-\frac{\mu_{h}\alpha_{k}}{4})z_{k}+\frac{9L^{2}\alpha_{k}}{\mu_{h}}y_{k}+\frac{9\alpha_{k}}{2\mu_{h}}\|\Delta f_{k}\|^{2},$
where again we use $\alpha_{k}\leq\frac{1}{2\mu_{h}}$ to simplify the terms.
$\blacksquare$
A.6 Proof of Lemma 5
By the update rule in (10),
$\omega_{k+1}-\omega^{\star}(\theta_{k+1})$ $=\omega_{k+1}-\omega^{\star}(\theta_{k+1})$
$=\left(\omega_{k}-\omega^{\star}(\theta_{k})\right)-\beta_{k}g_{k}-\left(\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\right)$
$=\left(\omega_{k}-\omega^{\star}(\theta_{k})\right)-\beta_{k}G(\theta_{k},\omega_{k})-\beta_{k}\Delta g_{k}-\left(\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\right).$
Taking the norm yields
$y_{k+1}$ $=\|\omega_{k+1}-\omega^{\star}(\theta_{k+1})\|^{2}$
$=\underbrace{\|\omega_{k}-\omega^{\star}(\theta_{k})-\beta_{k}G(\theta_{k},\omega_{k})\|^{2}}_{T_{1}}+\beta_{k}^{2}\|\Delta g_{k}\|^{2}+\underbrace{\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|^{2}}_{T_{2}}$
$\hskip 20.0pt\underbrace{-2\left(\omega_{k}-\omega^{\star}(\theta_{k})-\beta_{k}G(\theta_{k},\omega_{k})\right)^{\top}\left(\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\right)}_{T_{3}}$
$\hskip 20.0pt\underbrace{-2\beta_{k}\left(\omega_{k}-\omega^{\star}(\theta_{k})-\beta_{k}G(\theta_{k},\omega_{k})\right)^{\top}\Delta g_{k}}_{T_{4}}+\underbrace{2\beta_{k}\Delta g_{k}^{\top}\left(\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\right)}_{T_{5}}.$ (50)
We bound each term of (50) individually. First, since by definition $G(\theta,\omega^{\star}(\theta))=0$ for all $\theta$, we have
$T_{1}$ $=\|\omega_{k}-\omega^{\star}(\theta_{k})\|^{2}-2\beta_{k}\left(\omega_{k}-\omega^{\star}(\theta_{k})\right)^{\top}G(\theta_{k},\omega_{k})+\beta_{k}^{2}\|G(\theta_{k},\omega_{k})\|^{2}$
$=y_{k}-2\beta_{k}\left(\omega_{k}-\omega^{\star}(\theta_{k})\right)^{\top}G(\theta_{k},\omega_{k})+\beta_{k}^{2}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|^{2}$
$\leq y_{k}-2\mu_{G}\beta_{k}\|\omega_{k}-\omega^{\star}(\theta_{k})\|^{2}+L^{2}\beta_{k}^{2}\|\omega_{k}-\omega^{\star}(\theta_{k})\|^{2}$
$=(1-2\mu_{G}\beta_{k}+L^{2}\beta_{k}^{2})y_{k},$
where the first inequality follows from Assumption 3 and the Lipschitz continuity of $G$.
The term $T_{2}$ can be simply treated using the Lipschitz condition of $\omega^{\star}$
$T_{2}$ $=\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|^{2}\leq L^{2}\|\theta_{k+1}-\theta_{k}\|^{2}\leq L^{2}\alpha_{k}^{2}\left(\|\Delta f_{k}\|+L\sqrt{y_{k}}+L(L+1)\sqrt{z_{k}}\right)^{2}$
$\leq 3L^{2}\alpha_{k}^{2}\left(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+(L+1)^{4}z_{k}\right).$
By the Cauchy-Schwarz inequality,
$T_{3}$
$\leq 2\|\omega_{k}-\omega^{\star}(\theta_{k})-\beta_{k}G(\theta_{k},\omega_{k})\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|$
$\leq 2\|\omega_{k}-\omega^{\star}(\theta_{k})\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|+2\beta_{k}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|$
$\leq(2L+2L^{2}\beta_{k})\sqrt{y_{k}}\|\theta_{k+1}-\theta_{k}\|$
$\leq 4L\alpha_{k}\sqrt{y_{k}}\left(\|\Delta f_{k}\|+L\sqrt{y_{k}}+L(L+1)\sqrt{z_{k}}\right).$
where the second inequality follows from $G(\theta,\omega^{\star}(\theta))=0$ for all $\theta$, and the fourth inequality applies Lemma 1 and the step size rule $\beta_{k}\leq 1/L$.
We can bound $T_{4}$ following a similar line of analysis
$T_{4}$ $\leq 2\beta_{k}\|\omega_{k}-\omega^{\star}(\theta_{k})-\beta_{k}G(\theta_{k},\omega_{k})\|\|\Delta g_{k}\|\leq 4\beta_{k}\sqrt{y_{k}}\|\Delta g_{k}\|.$
Finally, we treat $T_{5}$ again by the Cauchy-Schwarz inequality
$T_{5}$ $\leq 2\beta_{k}\|\Delta g_{k}\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|$
$\leq 2L\alpha_{k}\beta_{k}\|\Delta g_{k}\|\Big{(}\|\Delta f_{k}\|+L\sqrt{y_{k}}+L(L+1)\sqrt{z_{k}}\Big{)},$
where the second inequality again uses Lemma 1.
Applying the bounds on $T_{1}$-$T_{5}$ to (50), we get
$y_{k+1}$ $\leq(1-2\mu_{G}\beta_{k}+L^{2}\beta_{k}^{2})y_{k}+\beta_{k}^{2}\|\Delta g_{k}\|^{2}+3L^{2}\alpha_{k}^{2}\left(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+(L+1)^{4}z_{k}\right)$
$\hskip 20.0pt+4L\alpha_{k}\sqrt{y_{k}}\left(\|\Delta f_{k}\|+L\sqrt{y_{k}}+L(L+1)\sqrt{z_{k}}\right)+4\beta_{k}\sqrt{y_{k}}\|\Delta g_{k}\|$
$\hskip 20.0pt+2L\alpha_{k}\beta_{k}\|\Delta g_{k}\|\Big{(}\|\Delta f_{k}\|+L\sqrt{y_{k}}+L(L+1)\sqrt{z_{k}}\Big{)}$
$\leq(1-2\mu_{G}\beta_{k}+L^{2}\beta_{k}^{2})y_{k}+\beta_{k}^{2}\|\Delta g_{k}\|^{2}+3L^{2}\alpha_{k}^{2}\left(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+(L+1)^{4}z_{k}\right)$
$\hskip 20.0pt+2L\alpha_{k}y_{k}+2L\alpha_{k}\|\Delta f_{k}\|^{2}+4L^{2}\alpha_{k}y_{k}+\frac{\mu_{G}}{4}\beta_{k}y_{k}+\frac{16(L+1)^{6}\alpha_{k}^{2}}{\mu_{G}\beta_{k}}z_{k}+\frac{\mu_{G}\beta_{k}}{2}y_{k}$
$\hskip 20.0pt+\frac{8\beta_{k}}{\mu_{G}}\|\Delta g_{k}\|^{2}+L\alpha_{k}\beta_{k}\|\Delta g_{k}\|^{2}+3L\alpha_{k}\beta_{k}\left(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+(L+1)^{4}z_{k}\right)$
$\leq(1-\mu_{G}\beta_{k})y_{k}-\left(\frac{\mu_{G}}{4}\beta_{k}-8L^{4}\alpha_{k}-L^{2}\beta_{k}^{2}\right)y_{k}$
$\hskip 20.0pt+\left(\frac{19(L+1)^{6}\alpha_{k}^{2}}{\mu_{G}\beta_{k}}+3(L+1)^{6}\alpha_{k}\beta_{k}\right)z_{k}+\left(\frac{8\beta_{k}}{\mu_{G}}+2L\beta_{k}^{2}\right)\|\Delta g_{k}\|^{2}+8L\alpha_{k}\|\Delta f_{k}\|^{2},$
where the second inequality follows from the fact that $2\|\boldsymbol{a}\|\|\boldsymbol{b}\|\leq c\|\boldsymbol{a}\|^{2}+\frac{1}{c}\|\boldsymbol{b}\|^{2}$ for any vector $\boldsymbol{a},\boldsymbol{b}$ and scalar $c>0$. We have also simplified terms using the conditions $\alpha_{k}\leq\beta_{k}\leq\frac{1}{72(L+1)^{6}}$ and $\beta_{k}\leq 1/\mu_{G}$.
$\blacksquare$
Appendix B Analysis for Functions Satisfying Polyak-Lojasiewicz Condition
The exact requirement on step sizes is
$\lambda_{k}=\frac{c_{\lambda}}{k+\tau+1},\quad\alpha_{k}=\frac{c_{\alpha}}{k+\tau+1},\quad\beta_{k}=\frac{c_{\beta}}{k+\tau+1},$ (51)
where $c_{\alpha},c_{\beta},c_{\lambda}$ need to be carefully balanced and $\tau$ selected sufficiently large such that $\alpha_{k}\leq\beta_{k}\leq\lambda_{k}\leq 1/4$, $c_{\beta}\leq 1/L$, $c_{\beta}\leq\frac{\mu_{G}}{320L^{4}}c_{\lambda}$, $c_{\beta}\leq\frac{\mu_{h}^{3}}{480(L+1)^{6}}$, $c_{\beta}\leq\frac{\mu_{G}}{8L^{2}}$, $\beta_{k}\leq\frac{\lambda_{k}}{4}\min\{(\frac{2L^{2}}{\mu_{h}^{3}}+58L^{2})^{-1},(50L^{2}+\frac{8}{\mu_{G}})^{-1}\}$, $\alpha_{k}\leq\frac{\mu_{G}\beta_{k}}{4\left(14L^{3}+\frac{132(L+1)^{6}}{\mu_{h}^{3}}\right)}$, $\alpha_{k}\leq\frac{\mu_{h}^{2}}{3072(L+1)^{6}}\lambda_{k}$, $\alpha_{k}\leq\beta_{k}$, and $c_{\alpha}\geq\frac{2}{\min\{\mu_{h},\mu_{G}\}}$.
The PL condition implies the following quadratic growth inequality, where $\text{Proj}_{\Theta^{\star}}(\theta)$ denotes the projection of $\theta$ to $\Theta^{\star}$ [Karimi et al., 2016].
$\frac{2}{\mu_{h}}\left(h(\theta)-h^{\star}\right)\geq\|\text{Proj}_{\Theta^{\star}}(\theta)-\theta\|^{2}.$ (52)
Equation (52) plays an important role in the proof of a lemma that we now introduce.
Lemma 6
Under Assumption 1-3 and Assumption 5, the sequence of variables $\{\theta_{k},\omega_{k},f_{k},g_{k}\}$ satisfy for all $k$
$\|f_{k}\|$ $\leq\|\Delta f_{k}\|+L\sqrt{y_{k}}+\frac{2L(L+1)}{\mu_{h}}\sqrt{x_{k}},$
$\|g_{k}\|$ $\leq\|\Delta g_{k}\|+L\sqrt{y_{k}},$
$\|\theta_{k+1}-\theta_{k}\|$ $\leq\alpha_{k}\left(\|\Delta f_{k}\|+L\sqrt{y_{k}}+\frac{2L(L+1)}{\mu_{h}}\sqrt{x_{k}}\right),$
$\|\omega_{k+1}-\omega_{k}\|$ $\leq\beta_{k}\left(\|\Delta g_{k}\|+L\sqrt{y_{k}}\right).$
Our analysis in this section also relies on a few more lemmas which we present below. The proofs of all lemmas can be found at the end of the section.
Lemma 7
Under Assumption 1-3 and Assumption 5 and the step sizes given in (51), we have the following bound on $\|\Delta f_{k+1}\|^{2}$
$\mathbb{E}[\|\Delta f_{k+1}\|^{2}]$ $\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+\frac{24(L+1)^{6}}{\mu_{h}}\alpha_{k}^{2}x_{k}\right]+2B\lambda_{k}^{2}.$
Lemma 8
Under Assumption 1-3 and Assumption 5 and the step sizes given in (51), we have the following bound on $\|\Delta g_{k+1}\|^{2}$
$\mathbb{E}[\|\Delta g_{k+1}\|^{2}]$ $\leq(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+\frac{24(L+1)^{6}}{\mu_{h}}\alpha_{k}^{2}x_{k}\right]+2B\lambda_{k}^{2}.$
Lemma 9
Under Assumption 1-3 and Assumption 5 and the step sizes given in (51), we have
$x_{k+1}$ $\leq(1-\mu_{h}\alpha_{k})x_{k}-(\frac{\mu_{h}}{4}\alpha_{k}-\frac{6(L+1)^{5}}{\mu_{h}^{2}}\alpha_{k}^{2})x_{k}$
$\hskip 20.0pt+(\frac{4L^{4}}{\mu_{h}^{3}}+2L)\alpha_{k}y_{k}+(\frac{2L^{2}\alpha_{k}}{\mu_{h}^{3}}+2L\alpha_{k}^{2})\|\Delta f_{k}\|^{2}.$
Lemma 10
Under Assumption 1-3 and Assumption 5 and the step sizes given in (51), we have
$y_{k+1}$ $\leq(1-\mu_{G}\beta_{k})y_{k}-\left(\frac{\mu_{G}}{2}\beta_{k}-\left(12L^{3}+\frac{128(L+1)^{6}}{\mu_{h}^{3}}\right)\alpha_{k}-L^{2}\beta_{k}^{2}\right)y_{k}$
$\hskip 20.0pt+\left(\frac{\mu_{h}}{8}\alpha_{k}+\frac{24(L+1)^{6}}{\mu_{h}^{2}}\alpha_{k}\beta_{k}\right)x_{k}+\left(\frac{8\beta_{k}}{\mu_{G}}+2L\beta_{k}^{2}\right)\|\Delta g_{k}\|^{2}+8L\alpha_{k}\|\Delta f_{k}\|^{2}.$
B.1 Proof of Theorem 2
Combining the results from Lemma 7-10 and simplifying the terms with the choice of step sizes given in (51),
$\mathbb{E}[V_{k+1}]$
$=\mathbb{E}[\|\Delta f_{k+1}\|^{2}+\|\Delta g_{k+1}\|^{2}+x_{k+1}+y_{k+1}]$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+\frac{24(L+1)^{6}}{\mu_{h}}\alpha_{k}^{2}x_{k}\right]+2B\lambda_{k}^{2}$
$\hskip 20.0pt+(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+\frac{24(L+1)^{6}}{\mu_{h}}\alpha_{k}^{2}x_{k}\right]+2B\lambda_{k}^{2}$
$\hskip 20.0pt+\mathbb{E}\Big{[}(1-\mu_{h}\alpha_{k})x_{k}-(\frac{\mu_{h}}{4}\alpha_{k}-\frac{6(L+1)^{5}}{\mu_{h}^{2}}\alpha_{k}^{2})x_{k}+(\frac{4L^{4}}{\mu_{h}^{3}}+2L)\alpha_{k}y_{k}+(\frac{2L^{2}\alpha_{k}}{\mu_{h}^{3}}+2L\alpha_{k}^{2})\|\Delta f_{k}\|^{2}$
$\hskip 20.0pt+(1-\mu_{G}\beta_{k})y_{k}-\left(\frac{\mu_{G}}{2}\beta_{k}-\left(12L^{3}+\frac{128(L+1)^{6}}{\mu_{h}^{3}}\right)\alpha_{k}-L^{2}\beta_{k}^{2}\right)y_{k}$
$\hskip 20.0pt+\left(\frac{\mu_{h}}{8}\alpha_{k}+\frac{24(L+1)^{6}}{\mu_{h}^{2}}\alpha_{k}\beta_{k}\right)x_{k}+\left(\frac{8\beta_{k}}{\mu_{G}}+2L\beta_{k}^{2}\right)\|\Delta g_{k}\|^{2}+8L\alpha_{k}\|\Delta f_{k}\|^{2}\Big{]}$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{48L^{2}\beta_{k}^{2}}{\lambda_{k}}-(\frac{2L^{2}}{\mu_{h}^{3}}+10L)\alpha_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-(50L^{2}+\frac{8}{\mu_{G}})\beta_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\frac{8}{\lambda_{k}}\mathbb{E}\left[10L^{4}\beta_{k}^{2}y_{k}+\frac{24(L+1)^{6}}{\mu_{h}}\alpha_{k}^{2}x_{k}\right]+4B\lambda_{k}^{2}$
$\hskip 20.0pt+(1-\mu_{h}\alpha_{k})\mathbb{E}[x_{k}]-(\frac{\mu_{h}}{4}\alpha_{k}-\frac{6(L+1)^{5}}{\mu_{h}^{2}}\alpha_{k}^{2})\mathbb{E}[x_{k}]+(\frac{4L^{4}}{\mu_{h}^{3}}+2L)\alpha_{k}\mathbb{E}[y_{k}]$
$\hskip 20.0pt+(1-\mu_{G}\beta_{k})\mathbb{E}[y_{k}]-\left(\frac{\mu_{G}}{2}\beta_{k}-\left(12L^{3}+\frac{128(L+1)^{6}}{\mu_{h}^{3}}\right)\alpha_{k}-L^{2}\beta_{k}^{2}\right)\mathbb{E}[y_{k}]$
$\hskip 20.0pt+\left(\frac{\mu_{h}}{8}\alpha_{k}+\frac{24(L+1)^{6}}{\mu_{h}^{2}}\alpha_{k}\beta_{k}\right)\mathbb{E}[x_{k}]$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]+(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]+(1-\mu_{h}\alpha_{k})\mathbb{E}[x_{k}]+(1-\mu_{G}\beta_{k})\mathbb{E}[y_{k}]+4B\lambda_{k}^{2}$
$\hskip 20.0pt-(\frac{\lambda_{k}}{4}-(\frac{2L^{2}}{\mu_{h}^{3}}+58L^{2})\beta_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-(50L^{2}+\frac{8}{\mu_{G}})\beta_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt-(\frac{\mu_{h}}{16}\alpha_{k}-\frac{30(L+1)^{6}}{\mu_{h}^{2}}\alpha_{k}\beta_{k})\mathbb{E}[x_{k}]-\left(\frac{\mu_{G}}{4}\beta_{k}-\left(14L^{3}+\frac{132(L+1)^{6}}{\mu_{h}^{3}}\right)\alpha_{k}-L^{2}\beta_{k}^{2}\right)\mathbb{E}[y_{k}],$
where we use the step size conditions $\beta_{k}\leq\frac{\mu_{G}}{320L^{4}}\lambda_{k}$ and $\alpha_{k}\leq\frac{\mu_{h}^{2}}{3072(L+1)^{6}}\lambda_{k}$.
Again, due to carefully selected step sizes, (specifically, $\beta_{k}\leq\frac{\mu_{h}^{3}}{480(L+1)^{6}}$, $\beta_{k}\leq\frac{\mu_{G}}{8L^{2}}$, $\beta_{k}\leq\frac{\lambda_{k}}{4}\min\{(\frac{2L^{2}}{\mu_{h}^{3}}+58L^{2})^{-1},(50L^{2}+\frac{8}{\mu_{G}})^{-1}\}$, and $\alpha_{k}\leq\frac{\mu_{G}\beta_{k}}{4\left(14L^{3}+\frac{132(L+1)^{6}}{\mu_{h}^{3}}\right)}$), we can make the last four terms of the inequality above non-positive. As a result, we have
$\mathbb{E}[V_{k+1}]$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]+(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]+(1-\mu_{h}\alpha_{k})\mathbb{E}[x_{k}]+(1-\mu_{G}\beta_{k})\mathbb{E}[y_{k}]+4B\lambda_{k}^{2}$
$\leq(1-\min\{\mu_{h},\mu_{G}\}\alpha_{k})\mathbb{E}[V_{k}]+4B\lambda_{k}^{2}$
$=(1-\frac{\min\{\mu_{h},\mu_{G}\}c_{\alpha}}{k+\tau+1})\mathbb{E}[V_{k}]+\frac{B}{4(k+1)^{2}}$
$\leq(1-\frac{2}{k+\tau+1})\mathbb{E}[V_{k}]+\frac{B}{4(k+1)^{2}}$
$\leq\frac{k+\tau-1}{k+\tau+1}\mathbb{E}[V_{k}]+\frac{B}{4(k+1)^{2}}.$
Multiplying by $(k+\tau+1)^{2}$ on both sides of the inequality, we have
$(k+\tau+1)^{2}\mathbb{E}[V_{k+1}]$ $\leq(k+\tau+1)(k+\tau-1)\mathbb{E}[V_{k}]+\frac{B(k+\tau+1)^{2}}{4(k+1)^{2}}$
$\leq(k+\tau)^{2}\mathbb{E}[V_{k}]+\left(\frac{B}{2}+\frac{B\tau^{2}}{2}\right)$
$\leq\tau^{2}V_{0}+\left(\frac{B}{2}+\frac{B\tau^{2}}{2}\right)(k+1).$
As a result,
$\mathbb{E}[V_{k+1}]$ $\leq\frac{B\tau^{2}(k+1)}{(k+\tau+1)^{2}}+\frac{\tau^{2}V_{0}}{(k+\tau+1)^{2}}$
$\leq\frac{B\tau^{2}}{k+\tau+1}+\frac{\tau^{2}V_{0}}{(k+\tau+1)^{2}},$
which obviously leads to the claimed result.
$\blacksquare$
B.2 Proof of Lemma 6
By definition,
$F(\theta,\omega^{\star}(\theta))=\nabla h(\theta)=0,\quad\forall\theta\in\Theta^{\star}$
which implies
$\|f_{k}\|$
$=\|\Delta f_{k}+F(\theta_{k},\omega_{k})-F(\theta_{k},\omega^{\star}(\theta_{k}))+F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\text{Proj}_{\Theta^{\star}}(\theta_{k}),\omega^{\star}(\text{Proj}_{\Theta^{\star}}(\theta_{k})))\|$
$\leq\|\Delta f_{k}\|+\|F(\theta_{k},\omega_{k})-F(\theta_{k},\omega^{\star}(\theta_{k}))\|+\|F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\text{Proj}_{\Theta^{\star}}(\theta_{k}),\omega^{\star}(\text{Proj}_{\Theta^{\star}}(\theta_{k})))\|$
$\leq\|\Delta f_{k}\|+L\|\omega_{k}-\omega^{\star}(\theta_{k})\|+L(L+1)\|\theta_{k}-\text{Proj}_{\Theta^{\star}}(\theta_{k})\|$
$\leq\|\Delta f_{k}\|+L\sqrt{y_{k}}+\frac{2L(L+1)}{\mu_{h}}\sqrt{x_{k}},$
where the second inequality follows from the Lipschitz continuity of $F$ and $\omega^{\star}$ and the final inequality applies (52).
Similarly, since $G(\theta,\omega^{\star}(\theta))=0$ for any $\theta$, we can derive
$\|g_{k}\|$ $=\|\Delta g_{k}+G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|$
$\leq\|\Delta g_{k}\|+\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|$
$\leq\|\Delta g_{k}\|+L\sqrt{y_{k}}.$
The bounds on $\|\theta_{k+1}-\theta_{k}\|$ and $\|\omega_{k+1}-\omega_{k}\|$ easily follow from the two inequalities above and (11).
$\blacksquare$
B.3 Proof of Lemma 7
By the update rule in (11), we have
$\Delta f_{k+1}$ $=f_{k+1}-F(\theta_{k+1},\omega_{k+1})$
$=(1-\lambda_{k})f_{k}+\lambda_{k}F(\theta_{k},\omega_{k},X_{k})-F(\theta_{k+1},\omega_{k+1})$
$=(1-\lambda_{k})f_{k}+\lambda_{k}F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})+\lambda_{k}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right)$
$=(1-\lambda_{k})\Delta f_{k}+F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})+\lambda_{k}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right).$
This leads to
$\|\Delta f_{k+1}\|^{2}$
$=(1-\lambda_{k})^{2}\|\Delta f_{k}\|^{2}+\Big{\|}\left(F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\right)+\lambda_{k}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right)\Big{\|}^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\Delta f_{k}^{\top}\left(F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\right)$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta f_{k}^{\top}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right)$
$\leq(1-\lambda_{k})^{2}\|\Delta f_{k}\|^{2}+2\|F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\|^{2}+2\lambda_{k}^{2}\|F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\|^{2}$
$\hskip 20.0pt+\frac{\lambda_{k}}{2}\|\Delta f_{k}\|^{2}+\frac{2}{\lambda_{k}}\left\|F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\right\|^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta f_{k}^{\top}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right)$
$\leq(1-\lambda_{k})\|\Delta f_{k}\|^{2}+2\lambda_{k}^{2}\|F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\|^{2}$
$\hskip 20.0pt+(\lambda_{k}^{2}-\frac{\lambda_{k}}{2})\|\Delta f_{k}\|^{2}+\frac{4}{\lambda_{k}}\left\|F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\right\|^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta f_{k}^{\top}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right),$ (53)
where the last inequality follows from $\lambda_{k}\leq 1$.
The second term on the right hand side of (53) is bounded in expectation, i.e.
$\mathbb{E}_{X_{k}\sim\xi}[\|F(\theta_{k},\omega_{k},X_{k})-F(\theta_{k},\omega_{k})\|^{2}]\leq B.$
In addition, the last term of (53) is zero in expectation since $\mathbb{E}[\Delta f_{k}^{\top}(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k}))\mid{\cal H}_{k-1}]=\Delta f_{k}^{\top}\mathbb{E}[(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k}))\mid{\cal H}_{k-1}]=0$. As a result, we have
$\mathbb{E}[\|\Delta f_{k+1}\|^{2}]$ $\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{2}-\lambda_{k}^{2})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}[\left\|F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\right\|^{2}]+2B\lambda_{k}^{2}.$ (54)
By Lemma 6 and the Lipschitz condition of $F$,
$\|F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\|^{2}$
$\leq 2L^{2}\|\theta_{k+1}-\theta_{k}\|^{2}+2L^{2}\|\omega_{k+1}-\omega_{k}\|^{2}$
$\leq 6L^{2}\alpha_{k}^{2}(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+\frac{4(L+1)^{4}}{\mu_{h}^{2}}x_{k})+4L^{2}\beta_{k}^{2}(\|\Delta g_{k}\|^{2}+L^{2}y_{k})$
$\leq 6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+\frac{24(L+1)^{6}}{\mu_{h}}\alpha_{k}^{2}x_{k}.$
Plugging this bound into (54), we get
$\mathbb{E}[\|\Delta f_{k+1}\|^{2}]$ $\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{2}-\lambda_{k}^{2})\mathbb{E}[\|\Delta f_{k}\|^{2}]+2B\lambda_{k}^{2}$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+\frac{24(L+1)^{6}}{\mu_{h}}\alpha_{k}^{2}x_{k}\right]$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+\frac{24(L+1)^{6}}{\mu_{h}}\alpha_{k}^{2}x_{k}\right]+2B\lambda_{k}^{2},$
where the last inequality follows from $\lambda_{k}\leq 1/4$.
$\blacksquare$
B.4 Proof of Lemma 8
By the update rule in (11), we have
$\Delta g_{k+1}$ $=g_{k+1}-G(\theta_{k+1},\omega_{k+1})$
$=(1-\lambda_{k})g_{k}+\lambda_{k}G(\theta_{k},\omega_{k},X_{k})-G(\theta_{k+1},\omega_{k+1})$
$=(1-\lambda_{k})g_{k}+\lambda_{k}G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})+\lambda_{k}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right)$
$=(1-\lambda_{k})\Delta g_{k}+G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})+\lambda_{k}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right).$
This leads to
$\|\Delta g_{k+1}\|^{2}$
$=(1-\lambda_{k})^{2}\|\Delta g_{k}\|^{2}+\Big{\|}\left(G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\right)+\lambda_{k}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right)\Big{\|}^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\Delta g_{k}^{\top}\left(G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\right)$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta g_{k}^{\top}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right)$
$\leq(1-\lambda_{k})^{2}\|\Delta g_{k}\|^{2}+2\|G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\|^{2}+2\lambda_{k}^{2}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\|^{2}$
$\hskip 20.0pt+\frac{\lambda_{k}}{2}\|\Delta g_{k}\|^{2}+\frac{2}{\lambda_{k}}\left\|G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\right\|^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta g_{k}^{\top}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right)$
$\leq(1-\lambda_{k})\|\Delta g_{k}\|^{2}+2\lambda_{k}^{2}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\|^{2}$
$\hskip 20.0pt+(\lambda_{k}^{2}-\frac{\lambda_{k}}{2})\|\Delta g_{k}\|^{2}+\frac{4}{\lambda_{k}}\left\|G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\right\|^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta g_{k}^{\top}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right).$ (55)
The second term on the right hand side of (55) is bounded in expectation, i.e.
$\mathbb{E}_{X_{k}\sim\xi}[\|G(\theta_{k},\omega_{k},X_{k})-G(\theta_{k},\omega_{k})\|^{2}]\leq B.$
In addition, the last term of (53) is zero in expectation since $\mathbb{E}[\Delta g_{k}^{\top}(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k}))\mid{\cal H}_{k-1}]=\Delta g_{k}^{\top}\mathbb{E}[(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k}))\mid{\cal H}_{k-1}]=0$. As a result, we have
$\mathbb{E}[\|\Delta g_{k+1}\|^{2}]$ $\leq(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{2}-\lambda_{k}^{2})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}[\left\|G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\right\|^{2}]+2B\lambda_{k}^{2}.$ (56)
By Lemma 6 and the Lipschitz condition of $G$,
$\|G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\|^{2}$
$\leq 2L^{2}\|\theta_{k+1}-\theta_{k}\|^{2}+2L^{2}\|\omega_{k+1}-\omega_{k}\|^{2}$
$\leq 6L^{2}\alpha_{k}^{2}(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+\frac{4(L+1)^{4}}{\mu_{h}^{2}}x_{k})+4L^{2}\beta_{k}^{2}(\|\Delta g_{k}\|^{2}+L^{2}y_{k})$
$\leq 6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+\frac{24(L+1)^{6}}{\mu_{h}}\alpha_{k}^{2}x_{k}.$
Plugging this bound into (56), we get
$\mathbb{E}[\|\Delta g_{k+1}\|^{2}]$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{2}-\lambda_{k}^{2})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+\frac{24(L+1)^{6}}{\mu_{h}}\alpha_{k}^{2}x_{k}\right]+2B\lambda_{k}^{2}$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+\frac{24(L+1)^{6}}{\mu_{h}}\alpha_{k}^{2}x_{k}\right]+2B\lambda_{k}^{2},$
where the last inequality follows from $\lambda_{k}\leq 1/4$.
$\blacksquare$
B.5 Proof of Lemma 9
From the update rule (10) and the Lipschitz gradient condition in Assumption 1, we have
$h(\theta_{k+1})$ $\leq h(\theta_{k})+\langle\nabla h(\theta_{k}),\theta_{k+1}-\theta_{k}\rangle+\frac{L}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$
$=h(\theta_{k})-\alpha_{k}\langle\nabla h(\theta_{k}),f_{k}\rangle+\frac{L\alpha_{k}^{2}}{2}\|f_{k}\|^{2}$
$=h(\theta_{k})-\alpha_{k}\langle\nabla h(\theta_{k}),\Delta f_{k}\rangle-\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega_{k})\rangle+\frac{L\alpha_{k}^{2}}{2}\|f_{k}\|^{2}$
$\leq h(\theta_{k})-\alpha_{k}\langle\nabla h(\theta_{k}),\Delta f_{k}\rangle-\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega^{\star}(\theta_{k}))\rangle$
$\hskip 20.0pt+\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\rangle$
$\hskip 20.0pt+\frac{3L\alpha_{k}^{2}}{2}\left(\|\Delta f_{k}\|^{2}+Ly_{k}+\frac{4(L+1)^{4}}{\mu_{h}^{2}}x_{k}\right)$
$=h(\theta_{k})-\alpha_{k}\|\nabla h(\theta_{k})\|^{2}-\alpha_{k}\langle\nabla h(\theta_{k}),\Delta f_{k}\rangle$
$\hskip 20.0pt+\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\rangle$
$\hskip 20.0pt+\frac{3L\alpha_{k}^{2}}{2}\left(\|\Delta f_{k}\|^{2}+Ly_{k}+\frac{4(L+1)^{4}}{\mu_{h}^{2}}x_{k}\right),$
where the second inequality applies Lemma 6.
As a result of Assumption 5, the inequality implies
$x_{x+1}$ $=h(\theta_{k+1})-h(\theta^{\star})$
$\leq h(\theta_{k})-h(\theta^{\star})-\alpha_{k}\|\nabla h(\theta_{k})\|^{2}-\alpha_{k}\langle\nabla h(\theta_{k}),\Delta f_{k}\rangle$
$\hskip 20.0pt+\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\rangle+\frac{3L\alpha_{k}^{2}}{2}\left(\|\Delta f_{k}\|^{2}+Ly_{k}+\frac{4(L+1)^{4}}{\mu_{h}^{2}}x_{k}\right)$
$\leq h(\theta_{k})-h(\theta^{\star})-2\mu_{h}\alpha_{k}\left(h(\theta_{k})-h(\theta^{\star})\right)-\alpha_{k}\langle\nabla h(\theta_{k}),\Delta f_{k}\rangle$
$\hskip 20.0pt+\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\rangle+\frac{3L\alpha_{k}^{2}}{2}\left(\|\Delta f_{k}\|^{2}+Ly_{k}+\frac{4(L+1)^{4}}{\mu_{h}^{2}}x_{k}\right)$
$\leq(1-2\mu_{h}\alpha_{k})x_{k}-\alpha_{k}\langle\nabla h(\theta_{k}),\Delta f_{k}\rangle$
$\hskip 20.0pt+\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\rangle+\frac{3L\alpha_{k}^{2}}{2}\left(\|\Delta f_{k}\|^{2}+Ly_{k}+\frac{4(L+1)^{4}}{\mu_{h}^{2}}x_{k}\right).$ (57)
To bound the second term on the right hand side of (57),
$-\alpha_{k}\langle\nabla h(\theta_{k}),\Delta f_{k}\rangle$ $\leq\alpha_{k}\|\nabla h(\theta_{k})\|\|\Delta f_{k}\|$
$=\alpha_{k}\|\nabla h(\theta_{k})-\nabla h(\text{Proj}_{\Theta^{\star}}(\theta_{k}))\|\|\Delta f_{k}\|$
$\leq L\alpha_{k}\|\theta_{k}-\text{Proj}_{\Theta^{\star}}(\theta_{k})\|\|\Delta f_{k}\|$
$\leq\frac{2L\alpha_{k}}{\mu_{h}}\sqrt{x_{k}}\|\Delta f_{k}\|$
$\leq\frac{\mu_{h}}{2}\alpha_{k}x_{k}+\frac{2L^{2}\alpha_{k}}{\mu_{h}^{3}}\|\Delta f_{k}\|^{2},$ (58)
where the third inequality plugs in the quadratic growth inequality (52) and the final inequality follows from the fact that $2\|\boldsymbol{a}\|\|\boldsymbol{b}\|\leq c\|\boldsymbol{a}\|^{2}+\frac{1}{c}\|\boldsymbol{b}\|^{2}$ for any vector $\boldsymbol{a},\boldsymbol{b}$ and scalar $c>0$.
The third term on the right hand side of (57) can be treated in a similar manner
$\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\rangle$ $\leq\alpha_{k}\|\nabla h(\theta_{k})\|\|F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\|$
$\leq\frac{2L\alpha_{k}}{\mu_{h}}\sqrt{x_{k}}\cdot\left(L\|\omega^{\star}(\theta_{k})-\omega_{k}\|\right)$
$=\frac{2L^{2}\alpha_{k}}{\mu_{h}}\sqrt{x_{k}}\sqrt{y_{k}}$
$\leq\frac{\mu_{h}}{4}\alpha_{k}x_{k}+\frac{4L^{4}\alpha_{k}}{\mu_{h}^{3}}y_{k}.$ (59)
Plugging (58) and (59) into (57), we get
$x_{k+1}$
$\leq(1-2\mu_{h}\alpha_{k})x_{k}-\alpha_{k}\langle\nabla h(\theta_{k}),\Delta f_{k}\rangle$
$\hskip 20.0pt+\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\rangle+\frac{3L\alpha_{k}^{2}}{2}\left(\|\Delta f_{k}\|^{2}+Ly_{k}+\frac{4(L+1)^{4}}{\mu_{h}^{2}}x_{k}\right)$
$\leq(1-2\mu_{h}\alpha_{k})x_{k}+\frac{\mu_{h}}{2}\alpha_{k}x_{k}+\frac{2L^{2}\alpha_{k}}{\mu_{h}^{3}}\|\Delta f_{k}\|^{2}$
$\hskip 20.0pt+\frac{\mu_{h}}{4}\alpha_{k}x_{k}+\frac{4L^{4}\alpha_{k}}{\mu_{h}^{3}}y_{k}+\frac{3L\alpha_{k}^{2}}{2}\left(\|\Delta f_{k}\|^{2}+Ly_{k}+\frac{4(L+1)^{4}}{\mu_{h}^{2}}x_{k}\right)$
$\leq(1-\mu_{h}\alpha_{k})x_{k}-(\frac{\mu_{h}}{4}\alpha_{k}-\frac{6(L+1)^{5}}{\mu_{h}^{2}}\alpha_{k}^{2})x_{k}+(\frac{4L^{4}}{\mu_{h}^{3}}+2L)\alpha_{k}y_{k}+(\frac{2L^{2}\alpha_{k}}{\mu_{h}^{3}}+2L\alpha_{k}^{2})\|\Delta f_{k}\|^{2}.$
$\blacksquare$
B.6 Proof of Lemma 10
By the update rule in (10),
$\omega_{k+1}-\omega^{\star}(\theta_{k+1})$ $=\omega_{k+1}-\omega^{\star}(\theta_{k+1})$
$=\left(\omega_{k}-\omega^{\star}(\theta_{k})\right)-\beta_{k}g_{k}-\left(\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\right)$
$=\left(\omega_{k}-\omega^{\star}(\theta_{k})\right)-\beta_{k}G(\theta_{k},\omega_{k})-\beta_{k}\Delta g_{k}-\left(\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\right).$
Taking the norm yields
$y_{k+1}$ $=\|\omega_{k+1}-\omega^{\star}(\theta_{k+1})\|^{2}$
$=\underbrace{\|\omega_{k}-\omega^{\star}(\theta_{k})-\beta_{k}G(\theta_{k},\omega_{k})\|^{2}}_{T_{1}}+\beta_{k}^{2}\|\Delta g_{k}\|^{2}+\underbrace{\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|^{2}}_{T_{2}}$
$\hskip 20.0pt\underbrace{-2\left(\omega_{k}-\omega^{\star}(\theta_{k})-\beta_{k}G(\theta_{k},\omega_{k})\right)^{\top}\left(\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\right)}_{T_{3}}$
$\hskip 20.0pt\underbrace{-2\beta_{k}\left(\omega_{k}-\omega^{\star}(\theta_{k})-\beta_{k}G(\theta_{k},\omega_{k})\right)^{\top}\Delta g_{k}}_{T_{4}}+\underbrace{2\beta_{k}\Delta g_{k}^{\top}\left(\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\right)}_{T_{5}}.$ (60)
We bound each term of (60) individually. First, since by definition $G(\theta,\omega^{\star}(\theta))=0$ for all $\theta$, we have
$T_{1}$ $=\|\omega_{k}-\omega^{\star}(\theta_{k})\|^{2}-2\beta_{k}\left(\omega_{k}-\omega^{\star}(\theta_{k})\right)^{\top}G(\theta_{k},\omega_{k})+\beta_{k}^{2}\|G(\theta_{k},\omega_{k})\|^{2}$
$=y_{k}-2\beta_{k}\left(\omega_{k}-\omega^{\star}(\theta_{k})\right)^{\top}G(\theta_{k},\omega_{k})+\beta_{k}^{2}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|^{2}$
$\leq y_{k}-2\mu_{G}\beta_{k}\|\omega_{k}-\omega^{\star}(\theta_{k})\|^{2}+L^{2}\beta_{k}^{2}\|\omega_{k}-\omega^{\star}(\theta_{k})\|^{2}$
$=(1-2\mu_{G}\beta_{k}+L^{2}\beta_{k}^{2})y_{k},$
where the first inequality follows from Assumption 3 and the Lipschitz continuity of $G$.
The term $T_{2}$ can be simply treated using the Lipschitz condition of $\omega^{\star}$
$T_{2}$ $=\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|^{2}\leq L^{2}\|\theta_{k+1}-\theta_{k}\|^{2}\leq L^{2}\alpha_{k}^{2}\left(\|\Delta f_{k}\|+L\sqrt{y_{k}}+\frac{2L(L+1)}{\mu_{h}}\sqrt{x_{k}}\right)^{2}$
$\leq 3L^{2}\alpha_{k}^{2}\left(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+\frac{4(L+1)^{4}}{\mu_{h}^{2}}x_{k}\right).$
By the Cauchy-Schwarz inequality,
$T_{3}$
$\leq 2\|\omega_{k}-\omega^{\star}(\theta_{k})-\beta_{k}G(\theta_{k},\omega_{k})\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|$
$\leq 2\|\omega_{k}-\omega^{\star}(\theta_{k})\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|+2\beta_{k}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|$
$\leq(2L+2L^{2}\beta_{k})\sqrt{y_{k}}\|\theta_{k+1}-\theta_{k}\|$
$\leq 4L\alpha_{k}\sqrt{y_{k}}\left(\|\Delta f_{k}\|+L\sqrt{y_{k}}+\frac{2L(L+1)}{\mu_{h}}\sqrt{x_{k}}\right).$
where the second inequality follows from $G(\theta,\omega^{\star}(\theta))=0$ for all $\theta$, and the fourth inequality applies Lemma 6 and the step size rule $\beta_{k}\leq 1/L$.
We can bound $T_{4}$ following a similar line of analysis
$T_{4}$ $\leq 2\beta_{k}\|\omega_{k}-\omega^{\star}(\theta_{k})-\beta_{k}G(\theta_{k},\omega_{k})\|\|\Delta g_{k}\|\leq 4\beta_{k}\sqrt{y_{k}}\|\Delta g_{k}\|.$
Finally, we treat $T_{5}$ again by the Cauchy-Schwarz inequality
$T_{5}$ $\leq 2\beta_{k}\|\Delta g_{k}\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|$
$\leq 2L\alpha_{k}\beta_{k}\|\Delta g_{k}\|\Big{(}\|\Delta f_{k}\|+L\sqrt{y_{k}}+\frac{2L(L+1)}{\mu_{h}}\sqrt{x_{k}}\Big{)},$
where the second inequality again uses Lemma 6.
Applying the bounds on $T_{1}$-$T_{5}$ to (60), we get
$y_{k+1}$ $\leq(1-2\mu_{G}\beta_{k}+L^{2}\beta_{k}^{2})y_{k}+\beta_{k}^{2}\|\Delta g_{k}\|^{2}+3L^{2}\alpha_{k}^{2}\left(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+\frac{4(L+1)^{4}}{\mu_{h}^{2}}x_{k}\right)$
$\hskip 20.0pt+4L\alpha_{k}\sqrt{y_{k}}\left(\|\Delta f_{k}\|+L\sqrt{y_{k}}+\frac{2L(L+1)}{\mu_{h}}\sqrt{x_{k}}\right)+4\beta_{k}\sqrt{y_{k}}\|\Delta g_{k}\|$
$\hskip 20.0pt+2L\alpha_{k}\beta_{k}\|\Delta g_{k}\|\Big{(}\|\Delta f_{k}\|+L\sqrt{y_{k}}+\frac{2L(L+1)}{\mu_{h}}\sqrt{x_{k}}\Big{)}$
$\leq(1-2\mu_{G}\beta_{k}+L^{2}\beta_{k}^{2})y_{k}+\beta_{k}^{2}\|\Delta g_{k}\|^{2}+3L^{2}\alpha_{k}^{2}\left(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+\frac{4(L+1)^{4}}{\mu_{h}^{2}}x_{k}\right)$
$\hskip 20.0pt+2L\alpha_{k}y_{k}+2L\alpha_{k}\|\Delta f_{k}\|^{2}+4L^{2}\alpha_{k}y_{k}+\frac{128(L+1)^{6}\alpha_{k}}{\mu_{h}^{3}}y_{k}+\frac{\mu_{h}\alpha_{k}}{8}x_{k}+\frac{\mu_{G}\beta_{k}}{2}y_{k}$
$\hskip 20.0pt+\frac{8\beta_{k}}{\mu_{G}}\|\Delta g_{k}\|^{2}+L\alpha_{k}\beta_{k}\|\Delta g_{k}\|^{2}+3L\alpha_{k}\beta_{k}\left(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+\frac{4(L+1)^{4}}{\mu_{h}^{2}}x_{k}\right)$
$\leq(1-\mu_{G}\beta_{k})y_{k}-\left(\frac{\mu_{G}}{2}\beta_{k}-\left(12L^{3}+\frac{128(L+1)^{6}}{\mu_{h}^{3}}\right)\alpha_{k}-L^{2}\beta_{k}^{2}\right)y_{k}$
$\hskip 20.0pt+\left(\frac{\mu_{h}}{8}\alpha_{k}+\frac{24(L+1)^{6}}{\mu_{h}^{2}}\alpha_{k}\beta_{k}\right)x_{k}+\left(\frac{8\beta_{k}}{\mu_{G}}+2L\beta_{k}^{2}\right)\|\Delta g_{k}\|^{2}+8L\alpha_{k}\|\Delta f_{k}\|^{2}$
where the second inequality follows from the fact that $2\|\boldsymbol{a}\|\|\boldsymbol{b}\|\leq c\|\boldsymbol{a}\|^{2}+\frac{1}{c}\|\boldsymbol{b}\|^{2}$ for any vector $\boldsymbol{a},\boldsymbol{b}$ and scalar $c>0$.
$\blacksquare$
Appendix C Analysis for Non-Convex Functions
Our analysis in this section relies on the following lemmas, the proofs of which can be found at the end of the section.
Lemma 11
Under Assumption 1-3, the sequence of variables $\{\theta_{k},\omega_{k},f_{k},g_{k}\}$ satisfy for all $k$
$\|f_{k}\|$ $\leq\|\Delta f_{k}\|+L\sqrt{y_{k}}+\|\nabla h(\theta_{k})\|,$
$\|g_{k}\|$ $\leq\|\Delta g_{k}\|+L\sqrt{y_{k}},$
$\|\theta_{k+1}-\theta_{k}\|$ $\leq\alpha_{k}\left(\|\Delta f_{k}\|+L\sqrt{y_{k}}+\|\nabla h(\theta_{k})\|\right),$
$\|\omega_{k+1}-\omega_{k}\|$ $\leq\beta_{k}\left(\|\Delta g_{k}\|+L\sqrt{y_{k}}\right).$
Lemma 12
Under Assumption 1-3 and the step sizes given in (18), we have the following bound on $\|\Delta f_{k+1}\|^{2}$
$\mathbb{E}[\|\Delta f_{k+1}\|^{2}]$ $\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6L^{2}\alpha_{k}^{2}\|\nabla h(\theta_{k})\|^{2}\right]+2B\lambda_{k}^{2}.$
Lemma 13
Under Assumption 1-3 and the step sizes given in (18), we have the following bound on $\|\Delta g_{k+1}\|^{2}$
$\mathbb{E}[\|\Delta g_{k+1}\|^{2}]$ $\leq(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6L^{2}\alpha_{k}^{2}\|\nabla h(\theta_{k})\|^{2}\right]+2B\lambda_{k}^{2}.$
Lemma 14
Under Assumption 1-3 and the step sizes in (18), we have
$h(\theta_{k+1})\leq h(\theta_{k})-\frac{\alpha_{k}}{4}\|\nabla h(\theta_{k})\|^{2}+\frac{5\alpha_{k}}{4}\|\Delta f_{k}\|^{2}+\frac{5L^{2}\alpha_{k}}{4}y_{k}.$
Lemma 15
Under Assumption 1-3 and the step sizes in (18), we have
$y_{k+1}$ $\leq(1-\mu_{G}\beta_{k})y_{k}-\left(\frac{\mu_{G}}{2}\beta_{k}-108L^{4}\alpha_{k}-L^{2}\beta_{k}^{2}\right)y_{k}$
$\hskip 20.0pt+\frac{\alpha_{k}}{8}\|\nabla h(\theta_{k})\|^{2}+\left(\frac{8\beta_{k}}{\mu_{G}}+2L\beta_{k}^{2}\right)\|\Delta g_{k}\|^{2}+8L\alpha_{k}\|\Delta f_{k}\|^{2}.$
C.1 Proof of Theorem 3
Combining the results from Lemma 12-15 and simplifying the terms with the choice of step sizes given in (18), we have
$\mathbb{E}[\|\Delta f_{k+1}\|^{2}+\|\Delta g_{k+1}\|^{2}+h(\theta_{k+1})+y_{k+1}]$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6L^{2}\alpha_{k}^{2}\|\nabla h(\theta_{k})\|^{2}\right]+2B\lambda_{k}^{2}$
$\hskip 20.0pt+(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6L^{2}\alpha_{k}^{2}\|\nabla h(\theta_{k})\|^{2}\right]+2B\lambda_{k}^{2}$
$\hskip 20.0pt+\mathbb{E}\Big{[}h(\theta_{k})-\frac{\alpha_{k}}{4}\|\nabla h(\theta_{k})\|^{2}+\frac{5\alpha_{k}}{4}\|\Delta f_{k}\|^{2}+\frac{5L^{2}\alpha_{k}}{4}y_{k}$
$\hskip 20.0pt+(1-\mu_{G}\beta_{k})y_{k}-\left(\frac{\mu_{G}}{2}\beta_{k}-108L^{4}\alpha_{k}-L^{2}\beta_{k}^{2}\right)y_{k}$
$\hskip 20.0pt+\frac{\alpha_{k}}{8}\|\nabla h(\theta_{k})\|^{2}+\left(\frac{8\beta_{k}}{\mu_{G}}+2L\beta_{k}^{2}\right)\|\Delta g_{k}\|^{2}+8L\alpha_{k}\|\Delta f_{k}\|^{2}\Big{]}$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{48L^{2}\beta_{k}^{2}}{\lambda_{k}}-\frac{37L}{4}\alpha_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-(50L^{2}+\frac{8}{\mu_{G}})\beta_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\mathbb{E}[h(\theta_{k})]-(\frac{\alpha_{k}}{8}-\frac{48L^{2}\alpha_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\nabla h(\theta_{k})\|^{2}]$
$\hskip 20.0pt+(1-\mu_{G}\beta_{k})\mathbb{E}[y_{k}]-\left(\frac{\mu_{G}}{2}\beta_{k}-110L^{4}\alpha_{k}-L^{2}\beta_{k}^{2}-\frac{80L^{4}\beta_{k}^{2}}{\lambda_{k}}\right)\mathbb{E}[y_{k}]+4B\lambda_{k}^{2}$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]+(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]+\mathbb{E}[h(\theta_{k})]-\frac{\alpha_{k}}{16}\mathbb{E}[\|\nabla h(\theta_{k})\|^{2}]$
$\hskip 20.0pt+(1-\mu_{G}\beta_{k})\mathbb{E}[y_{k}]+4B\lambda_{k}^{2},$
where we use the step size conditions $\beta_{k}\leq\frac{\lambda_{k}}{4}\min\{(\frac{37L}{4}+48L^{2})^{-1},(50L^{2}+\frac{8}{\mu_{G}})^{-1}\}$, $\alpha_{k}\leq\frac{1}{768L^{2}}\lambda_{k}$, $\beta_{k}\leq\frac{\mu_{G}}{320L^{4}}\lambda_{k}$, $\beta_{k}\leq\frac{\mu_{G}}{8L^{2}}$, and $\alpha_{k}\leq\frac{\mu_{G}}{880L^{4}}\beta_{k}$.
This inequality obviously implies
$\alpha_{k}\mathbb{E}[\|\nabla h(\theta_{k})\|^{2}]$ $\leq 16\mathbb{E}[\|\Delta f_{k}\|^{2}-\|\Delta f_{k+1}\|^{2}]+16\mathbb{E}[\|\Delta g_{k}\|^{2}-\|\Delta g_{k+1}\|^{2}]$
$\hskip 20.0pt+16\mathbb{E}[h(\theta_{k})-h(\theta_{k+1})]+16\mathbb{E}[y_{k}-y_{k+1}]+64B\lambda_{k}^{2}.$
Since $\|\Delta f_{k}\|^{2}$, $\|\Delta g_{k}\|^{2}$, and $y_{k}$ are non-negative and $h(\theta_{k})\geq h(\theta^{\star})$, we have the following inequality by summing up over $k$
$\sum_{t=0}^{k}\alpha_{t}\mathbb{E}[\|\nabla h(\theta_{t})\|^{2}]$ $\leq 16\left(\|\Delta f_{0}\|^{2}+\|\Delta g_{0}\|^{2}+(h(\theta_{0})-h(\theta^{\star}))+y_{0}\right)+64B\sum_{t=0}^{k}\lambda_{t}^{2}$
It is a standard result that (see, for example, Lemma 3 of Zeng et al. [2024])
$\sum_{t=0}^{k}\alpha_{t}=\alpha_{0}\sum_{t=0}^{k}\frac{1}{(t+1)^{1/2}}\geq\frac{\alpha_{0}(k+1)^{1/2}}{2},$
$\sum_{t=0}^{k}\lambda_{t}^{2}=\lambda_{0}^{2}\sum_{t=0}^{k}\frac{1}{t+1}\leq 2\lambda_{0}^{2}\log(k+2)$
Finally,
$\min_{t\leq k}\mathbb{E}\left[\|\nabla h(\theta_{t})\|^{2}\right]$
$\leq\frac{1}{\sum_{t^{\prime}=0}^{k}\alpha_{t^{\prime}}}\sum_{t=0}^{k}\alpha_{t}\mathbb{E}\left[\|\nabla h(\theta_{t})\|^{2}\right]$
$\leq\frac{32}{\alpha_{0}(k+1)^{1/2}}\left(\|\Delta f_{0}\|^{2}+\|\Delta g_{0}\|^{2}+(h(\theta_{0})-h(\theta^{\star}))+y_{0}\right)+\frac{4\log(k+2)}{\alpha_{0}(k+1)^{1/2}}.$
$\blacksquare$
C.2 Proof of Lemma 11
By the update rule,
$\|f_{k}\|$ $=\|\Delta f_{k}+F(\theta_{k},\omega_{k})-F(\theta_{k},\omega^{\star}(\theta_{k}))+F(\theta_{k},\omega^{\star}(\theta_{k}))\|$
$\leq\|\Delta f_{k}\|+\|F(\theta_{k},\omega_{k})-F(\theta_{k},\omega^{\star}(\theta_{k}))\|+\|F(\theta_{k},\omega^{\star}(\theta_{k}))\|$
$=\|\Delta f_{k}\|+\|F(\theta_{k},\omega_{k})-F(\theta_{k},\omega^{\star}(\theta_{k}))\|+\|\nabla h(\theta_{k})\|$
$\leq\|\Delta f_{k}\|+L\|\omega_{k}-\omega^{\star}(\theta_{k})\|+\|\nabla h(\theta_{k})\|$
$\leq\|\Delta f_{k}\|+L\sqrt{y_{k}}+\|\nabla h(\theta_{k})\|,$
where the second inequality follows from the Lipschitz continuity of $F$.
Similarly, since $G(\theta,\omega^{\star}(\theta))=0$ for any $\theta$, we can derive
$\|g_{k}\|$ $=\|\Delta g_{k}+G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|$
$\leq\|\Delta g_{k}\|+\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|$
$\leq\|\Delta g_{k}\|+L\sqrt{y_{k}}.$
The bounds on $\|\theta_{k+1}-\theta_{k}\|$ and $\|\omega_{k+1}-\omega_{k}\|$ easily follow from the two inequalities above and (11).
$\blacksquare$
C.3 Proof of Lemma 12
By the update rule in (11), we have
$\Delta f_{k+1}$ $=f_{k+1}-F(\theta_{k+1},\omega_{k+1})$
$=(1-\lambda_{k})f_{k}+\lambda_{k}F(\theta_{k},\omega_{k},X_{k})-F(\theta_{k+1},\omega_{k+1})$
$=(1-\lambda_{k})f_{k}+\lambda_{k}F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})+\lambda_{k}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right)$
$=(1-\lambda_{k})\Delta f_{k}+F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})+\lambda_{k}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right).$
This leads to
$\|\Delta f_{k+1}\|^{2}$
$=(1-\lambda_{k})^{2}\|\Delta f_{k}\|^{2}+\Big{\|}\left(F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\right)+\lambda_{k}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right)\Big{\|}^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\Delta f_{k}^{\top}\left(F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\right)$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta f_{k}^{\top}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right)$
$\leq(1-\lambda_{k})^{2}\|\Delta f_{k}\|^{2}+2\|F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\|^{2}+2\lambda_{k}^{2}\|F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\|^{2}$
$\hskip 20.0pt+\frac{\lambda_{k}}{2}\|\Delta f_{k}\|^{2}+\frac{2}{\lambda_{k}}\left\|F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\right\|^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta f_{k}^{\top}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right)$
$\leq(1-\lambda_{k})\|\Delta f_{k}\|^{2}+2\lambda_{k}^{2}\|F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\|^{2}$
$\hskip 20.0pt+(\lambda_{k}^{2}-\frac{\lambda_{k}}{2})\|\Delta f_{k}\|^{2}+\frac{4}{\lambda_{k}}\left\|F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\right\|^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta f_{k}^{\top}\left(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k})\right),$ (61)
where the last inequality follows from $\lambda_{k}\leq 1$.
The second term on the right hand side of (61) is bounded in expectation, i.e.
$\mathbb{E}_{X_{k}\sim\xi}[\|F(\theta_{k},\omega_{k},X_{k})-F(\theta_{k},\omega_{k})\|^{2}]\leq B.$
In addition, the last term of (61) is zero in expectation since $\mathbb{E}[\Delta f_{k}^{\top}(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k}))\mid{\cal H}_{k-1}]=\Delta f_{k}^{\top}\mathbb{E}[(F(\theta_{k},\omega_{k})-F(\theta_{k},\omega_{k},X_{k}))\mid{\cal H}_{k-1}]=0$. As a result, we have
$\mathbb{E}[\|\Delta f_{k+1}\|^{2}]$ $\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{2}-\lambda_{k}^{2})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}[\left\|F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\right\|^{2}]+2B\lambda_{k}^{2}.$ (62)
By Lemma 11 and the Lipschitz condition of $F$,
$\|F(\theta_{k},\omega_{k})-F(\theta_{k+1},\omega_{k+1})\|^{2}$
$\leq 2L^{2}\|\theta_{k+1}-\theta_{k}\|^{2}+2L^{2}\|\omega_{k+1}-\omega_{k}\|^{2}$
$\leq 6L^{2}\alpha_{k}^{2}(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+\|\nabla h(\theta_{k})\|^{2})+4L^{2}\beta_{k}^{2}(\|\Delta g_{k}\|^{2}+L^{2}y_{k})$
$\leq 6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6L^{2}\alpha_{k}^{2}\|\nabla h(\theta_{k})\|^{2}.$
Plugging this bound into (62), we get
$\mathbb{E}[\|\Delta f_{k+1}\|^{2}]$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{2}-\lambda_{k}^{2})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6L^{2}\alpha_{k}^{2}\|\nabla h(\theta_{k})\|^{2}\right]+2B\lambda_{k}^{2}$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta f_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta f_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6L^{2}\alpha_{k}^{2}\|\nabla h(\theta_{k})\|^{2}\right]+2B\lambda_{k}^{2},$
where the last inequality follows from $\lambda_{k}\leq 1/4$.
$\blacksquare$
C.4 Proof of Lemma 13
By the update rule in (11), we have
$\Delta g_{k+1}$ $=g_{k+1}-G(\theta_{k+1},\omega_{k+1})$
$=(1-\lambda_{k})g_{k}+\lambda_{k}G(\theta_{k},\omega_{k},X_{k})-G(\theta_{k+1},\omega_{k+1})$
$=(1-\lambda_{k})g_{k}+\lambda_{k}G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})+\lambda_{k}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right)$
$=(1-\lambda_{k})\Delta g_{k}+G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})+\lambda_{k}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right).$
This leads to
$\|\Delta g_{k+1}\|^{2}$
$=(1-\lambda_{k})^{2}\|\Delta g_{k}\|^{2}+\Big{\|}\left(G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\right)+\lambda_{k}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right)\Big{\|}^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\Delta g_{k}^{\top}\left(G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\right)$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta g_{k}^{\top}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right)$
$\leq(1-\lambda_{k})^{2}\|\Delta g_{k}\|^{2}+2\|G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\|^{2}+2\lambda_{k}^{2}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\|^{2}$
$\hskip 20.0pt+\frac{\lambda_{k}}{2}\|\Delta g_{k}\|^{2}+\frac{2}{\lambda_{k}}\left\|G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\right\|^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta g_{k}^{\top}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right)$
$\leq(1-\lambda_{k})\|\Delta g_{k}\|^{2}+2\lambda_{k}^{2}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\|^{2}$
$\hskip 20.0pt+(\lambda_{k}^{2}-\frac{\lambda_{k}}{2})\|\Delta g_{k}\|^{2}+\frac{4}{\lambda_{k}}\left\|G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\right\|^{2}$
$\hskip 20.0pt+2(1-\lambda_{k})\lambda_{k}\Delta g_{k}^{\top}\left(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k})\right).$ (63)
The second term on the right hand side of (63) is bounded in expectation, i.e.
$\mathbb{E}_{X_{k}\sim\xi}[\|G(\theta_{k},\omega_{k},X_{k})-G(\theta_{k},\omega_{k})\|^{2}]\leq B.$
In addition, the last term of (61) is zero in expectation since $\mathbb{E}[\Delta g_{k}^{\top}(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k}))\mid{\cal H}_{k-1}]=\Delta g_{k}^{\top}\mathbb{E}[(G(\theta_{k},\omega_{k})-G(\theta_{k},\omega_{k},X_{k}))\mid{\cal H}_{k-1}]=0$. As a result, we have
$\mathbb{E}[\|\Delta g_{k+1}\|^{2}]$ $\leq(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{2}-\lambda_{k}^{2})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}[\left\|G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\right\|^{2}]+2B\lambda_{k}^{2}.$ (64)
By Lemma 11 and the Lipschitz condition of $G$,
$\|G(\theta_{k},\omega_{k})-G(\theta_{k+1},\omega_{k+1})\|^{2}$
$\leq 2L^{2}\|\theta_{k+1}-\theta_{k}\|^{2}+2L^{2}\|\omega_{k+1}-\omega_{k}\|^{2}$
$\leq 6L^{2}\alpha_{k}^{2}(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+\|\nabla h(\theta_{k})\|^{2})+4L^{2}\beta_{k}^{2}(\|\Delta g_{k}\|^{2}+L^{2}y_{k})$
$\leq 6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6L^{2}\alpha_{k}^{2}\|\nabla h(\theta_{k})\|^{2}.$
Plugging this bound into (64), we get
$\mathbb{E}[\|\Delta g_{k+1}\|^{2}]$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{2}-\lambda_{k}^{2})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+6L^{2}\beta_{k}^{2}\|\Delta g_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6L^{2}\alpha_{k}^{2}\|\nabla h(\theta_{k})\|^{2}\right]+2B\lambda_{k}^{2}$
$\leq(1-\lambda_{k})\mathbb{E}[\|\Delta g_{k}\|^{2}]-(\frac{\lambda_{k}}{4}-\frac{24L^{2}\beta_{k}^{2}}{\lambda_{k}})\mathbb{E}[\|\Delta g_{k}\|^{2}]$
$\hskip 20.0pt+\frac{4}{\lambda_{k}}\mathbb{E}\left[6L^{2}\beta_{k}^{2}\|\Delta f_{k}\|^{2}+10L^{4}\beta_{k}^{2}y_{k}+6L^{2}\alpha_{k}^{2}\|\nabla h(\theta_{k})\|^{2}\right]+2B\lambda_{k}^{2},$
where the last inequality follows from $\lambda_{k}\leq 1/4$.
$\blacksquare$
C.5 Proof of Lemma 14
From the update rule (10) and the Lipschitz gradient condition in Assumption 1, we have
$h(\theta_{k+1})$ $\leq h(\theta_{k})+\langle\nabla h(\theta_{k}),\theta_{k+1}-\theta_{k}\rangle+\frac{L}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$
$=h(\theta_{k})-\alpha_{k}\langle\nabla h(\theta_{k}),f_{k}\rangle+\frac{L\alpha_{k}^{2}}{2}\|f_{k}\|^{2}$
$=h(\theta_{k})-\alpha_{k}\langle\nabla h(\theta_{k}),\Delta f_{k}\rangle-\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega^{\star}(\theta_{k}))\rangle$
$\hskip 20.0pt+\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\rangle+\frac{L\alpha_{k}^{2}}{2}\|f_{k}\|^{2}$
$=h(\theta_{k})-\alpha_{k}\|\nabla h(\theta_{k})\|^{2}-\alpha_{k}\langle\nabla h(\theta_{k}),\Delta f_{k}\rangle$
$\hskip 20.0pt+\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\rangle+\frac{L\alpha_{k}^{2}}{2}\|f_{k}\|^{2}.$ (65)
To bound the second term on the right hand side of (65),
$-\alpha_{k}\langle\nabla h(\theta_{k}),\Delta f_{k}\rangle$ $\leq\alpha_{k}\|\nabla h(\theta_{k})\|\|\Delta f_{k}\|\leq\frac{\alpha_{k}}{4}\|\nabla h(\theta_{k})\|^{2}+\alpha_{k}\|\Delta f_{k}\|^{2},$ (66)
where the final inequality follows from the fact that $2\|\boldsymbol{a}\|\|\boldsymbol{b}\|\leq c\|\boldsymbol{a}\|^{2}+\frac{1}{c}\|\boldsymbol{b}\|^{2}$ for any vector $\boldsymbol{a},\boldsymbol{b}$ and scalar $c>0$.
The third term on the right hand side of (65) can be treated in a similar manner
$\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\rangle$ $\leq\alpha_{k}\|\nabla h(\theta_{k})\|\|F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\|$
$\leq L\alpha_{k}\|\nabla h(\theta_{k})\|\|\omega^{\star}(\theta_{k})-\omega_{k}\|$
$\leq\frac{\alpha_{k}}{4}\|\nabla h(\theta_{k})\|^{2}+L^{2}\alpha_{k}y_{k}.$ (67)
Plugging (66) and (67) into (65), we get
$h(\theta_{k+1})$
$\leq h(\theta_{k})-\alpha_{k}\|\nabla h(\theta_{k})\|^{2}-\alpha_{k}\langle\nabla h(\theta_{k}),\Delta f_{k}\rangle$
$\hskip 20.0pt+\alpha_{k}\langle\nabla h(\theta_{k}),F(\theta_{k},\omega^{\star}(\theta_{k}))-F(\theta_{k},\omega_{k})\rangle+\frac{L\alpha_{k}^{2}}{2}\|f_{k}\|^{2}$
$\leq h(\theta_{k})-\alpha_{k}\|\nabla h(\theta_{k})\|^{2}+\frac{\alpha_{k}}{4}\|\nabla h(\theta_{k})\|^{2}+\alpha_{k}\|\Delta f_{k}\|^{2}+\frac{\alpha_{k}}{4}\|\nabla h(\theta_{k})\|^{2}$
$\hskip 20.0pt+L^{2}\alpha_{k}y_{k}+\frac{L\alpha_{k}^{2}}{2}\|f_{k}\|^{2}$
$\leq h(\theta_{k})-\frac{\alpha_{k}}{2}\|\nabla h(\theta_{k})\|^{2}+\alpha_{k}\|\Delta f_{k}\|^{2}+L^{2}\alpha_{k}y_{k}+\frac{3L\alpha_{k}^{2}}{2}\left(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+\|\nabla h(\theta_{k})\|^{2}\right)$
$\leq h(\theta_{k})-\frac{\alpha_{k}}{4}\|\nabla h(\theta_{k})\|^{2}+\frac{5\alpha_{k}}{4}\|\Delta f_{k}\|^{2}+\frac{5L^{2}\alpha_{k}}{4}y_{k},$
where the third inequality plugs in the result of Lemma 11, and the final inequality uses $\alpha_{k}\leq\frac{1}{6L}$.
$\blacksquare$
C.6 Proof of Lemma 15
By the update rule in (10),
$\omega_{k+1}-\omega^{\star}(\theta_{k+1})$ $=\omega_{k+1}-\omega^{\star}(\theta_{k+1})$
$=\left(\omega_{k}-\omega^{\star}(\theta_{k})\right)-\beta_{k}g_{k}-\left(\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\right)$
$=\left(\omega_{k}-\omega^{\star}(\theta_{k})\right)-\beta_{k}G(\theta_{k},\omega_{k})-\beta_{k}\Delta g_{k}-\left(\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\right).$
Taking the norm yields
$y_{k+1}$ $=\|\omega_{k+1}-\omega^{\star}(\theta_{k+1})\|^{2}$
$=\underbrace{\|\omega_{k}-\omega^{\star}(\theta_{k})-\beta_{k}G(\theta_{k},\omega_{k})\|^{2}}_{T_{1}}+\beta_{k}^{2}\|\Delta g_{k}\|^{2}+\underbrace{\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|^{2}}_{T_{2}}$
$\hskip 20.0pt\underbrace{-2\left(\omega_{k}-\omega^{\star}(\theta_{k})-\beta_{k}G(\theta_{k},\omega_{k})\right)^{\top}\left(\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\right)}_{T_{3}}$
$\hskip 20.0pt\underbrace{-2\beta_{k}\left(\omega_{k}-\omega^{\star}(\theta_{k})-\beta_{k}G(\theta_{k},\omega_{k})\right)^{\top}\Delta g_{k}}_{T_{4}}+\underbrace{2\beta_{k}\Delta g_{k}^{\top}\left(\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\right)}_{T_{5}}.$ (68)
We bound each term of (68) individually. First, since by definition $G(\theta,\omega^{\star}(\theta))=0$ for all $\theta$, we have
$T_{1}$ $=\|\omega_{k}-\omega^{\star}(\theta_{k})\|^{2}-2\beta_{k}\left(\omega_{k}-\omega^{\star}(\theta_{k})\right)^{\top}G(\theta_{k},\omega_{k})+\beta_{k}^{2}\|G(\theta_{k},\omega_{k})\|^{2}$
$=y_{k}-2\beta_{k}\left(\omega_{k}-\omega^{\star}(\theta_{k})\right)^{\top}G(\theta_{k},\omega_{k})+\beta_{k}^{2}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|^{2}$
$\leq y_{k}-2\mu_{G}\beta_{k}\|\omega_{k}-\omega^{\star}(\theta_{k})\|^{2}+L^{2}\beta_{k}^{2}\|\omega_{k}-\omega^{\star}(\theta_{k})\|^{2}$
$=(1-2\mu_{G}\beta_{k}+L^{2}\beta_{k}^{2})y_{k},$
where the first inequality follows from Assumption 3 and the Lipschitz continuity of $G$.
The term $T_{2}$ can be simply treated using the Lipschitz condition of $\omega^{\star}$
$T_{2}$ $=\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|^{2}\leq L^{2}\|\theta_{k+1}-\theta_{k}\|^{2}\leq L^{2}\alpha_{k}^{2}\left(\|\Delta f_{k}\|+L\sqrt{y_{k}}+\|\nabla h(\theta_{k})\|\right)^{2}$
$\leq 3L^{2}\alpha_{k}^{2}\left(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+\|\nabla h(\theta_{k})\|^{2}\right).$
By the Cauchy-Schwarz inequality,
$T_{3}$
$\leq 2\|\omega_{k}-\omega^{\star}(\theta_{k})-\beta_{k}G(\theta_{k},\omega_{k})\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|$
$\leq 2\|\omega_{k}-\omega^{\star}(\theta_{k})\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|+2\beta_{k}\|G(\theta_{k},\omega_{k})-G(\theta_{k},\omega^{\star}(\theta_{k}))\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|$
$\leq(2L+2L^{2}\beta_{k})\sqrt{y_{k}}\|\theta_{k+1}-\theta_{k}\|$
$\leq 4L\alpha_{k}\sqrt{y_{k}}\left(\|\Delta f_{k}\|+L\sqrt{y_{k}}+\|\nabla h(\theta_{k})\|\right).$
where the second inequality follows from $G(\theta,\omega^{\star}(\theta))=0$ for all $\theta$, and the fourth inequality applies Lemma 11 and the step size rule $\beta_{k}\leq 1/L$.
We can bound $T_{4}$ following a similar line of analysis
$T_{4}$ $\leq 2\beta_{k}\|\omega_{k}-\omega^{\star}(\theta_{k})-\beta_{k}G(\theta_{k},\omega_{k})\|\|\Delta g_{k}\|\leq 4\beta_{k}\sqrt{y_{k}}\|\Delta g_{k}\|.$
Finally, we treat $T_{5}$ again by the Cauchy-Schwarz inequality
$T_{5}$ $\leq 2\beta_{k}\|\Delta g_{k}\|\|\omega^{\star}(\theta_{k+1})-\omega^{\star}(\theta_{k})\|$
$\leq 2L\alpha_{k}\beta_{k}\|\Delta g_{k}\|\Big{(}\|\Delta f_{k}\|+L\sqrt{y_{k}}+\|\nabla h(\theta_{k})\|\Big{)},$
where the second inequality again uses Lemma 11.
Applying the bounds on $T_{1}$-$T_{5}$ to (68), we get
$y_{k+1}$
$\leq(1-2\mu_{G}\beta_{k}+L^{2}\beta_{k}^{2})y_{k}+\beta_{k}^{2}\|\Delta g_{k}\|^{2}+3L^{2}\alpha_{k}^{2}\left(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+\|\nabla h(\theta_{k})\|^{2}\right)$
$\hskip 20.0pt+4L\alpha_{k}\sqrt{y_{k}}\left(\|\Delta f_{k}\|+L\sqrt{y_{k}}+\|\nabla h(\theta_{k})\|\right)+4\beta_{k}\sqrt{y_{k}}\|\Delta g_{k}\|$
$\hskip 20.0pt+2L\alpha_{k}\beta_{k}\|\Delta g_{k}\|\Big{(}\|\Delta f_{k}\|+L\sqrt{y_{k}}+\|\nabla h(\theta_{k})\|\Big{)}$
$\leq(1-2\mu_{G}\beta_{k}+L^{2}\beta_{k}^{2})y_{k}+\beta_{k}^{2}\|\Delta g_{k}\|^{2}+3L^{2}\alpha_{k}^{2}\left(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+\|\nabla h(\theta_{k})\|^{2}\right)$
$\hskip 20.0pt+2L\alpha_{k}y_{k}+2L\alpha_{k}\|\Delta f_{k}\|^{2}+4L^{2}\alpha_{k}y_{k}+96L^{2}\alpha_{k}y_{k}+\frac{\alpha_{k}}{24}\|\nabla h(\theta_{k})\|^{2}+\frac{\mu_{G}\beta_{k}}{2}y_{k}+\frac{8\beta_{k}}{\mu_{G}}\|\Delta g_{k}\|^{2}$
$\hskip 20.0pt+L\alpha_{k}\beta_{k}\|\Delta g_{k}\|^{2}+3L\alpha_{k}\beta_{k}\left(\|\Delta f_{k}\|^{2}+L^{2}y_{k}+\|\nabla h(\theta_{k})\|^{2}\right)$
$\leq(1-\mu_{G}\beta_{k})y_{k}-\left(\frac{\mu_{G}}{2}\beta_{k}-108L^{4}\alpha_{k}-L^{2}\beta_{k}^{2}\right)y_{k}$
$\hskip 20.0pt+\frac{\alpha_{k}}{8}\|\nabla h(\theta_{k})\|^{2}+\left(\frac{8\beta_{k}}{\mu_{G}}+2L\beta_{k}^{2}\right)\|\Delta g_{k}\|^{2}+8L\alpha_{k}\|\Delta f_{k}\|^{2},$
where the second inequality follows from the fact that $2\|\boldsymbol{a}\|\|\boldsymbol{b}\|\leq c\|\boldsymbol{a}\|^{2}+\frac{1}{c}\|\boldsymbol{b}\|^{2}$ for any vector $\boldsymbol{a},\boldsymbol{b}$ and scalar $c>0$. We have also simplified terms using the conditions $\alpha_{k}\leq\frac{1}{72L^{2}}$ and $\beta_{k}\leq\frac{1}{72L}$.
$\blacksquare$
Appendix D Simulations Details
We generate completely random MDPs with transition probability matrix ${\cal P}\in\mathbb{R}^{|{\cal S}|\times|{\cal A}|\times|{\cal S}|}$ drawn i.i.d. from $\text{Unif}(0,1)$ and then normalized such that
$\sum_{s^{\prime}\in{\cal S}}{\cal P}(s^{\prime}\mid s,a)=1,\quad\forall s,a.$
The behavior policy $\pi_{b}$ is chosen to be uniform for all states. The policy $\pi$ to be evaluated is generated randomly by first drawing $\psi\in\mathbb{R}^{|{\cal S}|\times|{\cal A}|}$ such that $\psi_{s,a}\sim N(0,1)$, which produces the policy $\pi$ through the softmax function
$\pi(a\mid s)=\frac{\exp(\psi_{s,a})}{\sum_{a^{\prime}}\exp(\psi_{s,a^{\prime}})},\quad\forall s,a.$
We draw the feature matrix $\Phi\in\mathbb{R}^{|{\cal S}|\times dx}$ entry-wise i.i.d. from $N(0,1)$. The value function $V\in\mathbb{R}^{|{\cal S}|}$ is the solution to the Bellman equation
$V=R+\gamma P^{\pi}V,$
where the matrix $P^{\pi}\in\mathbb{R}^{|{\cal S}|\times|{\cal S}|}$ is the state transition matrix under policy $\pi$
$P_{s,s^{\prime}}=\sum_{a}{\cal P}(s^{\prime}\mid s,a)\pi(a\mid s)$
We would like to design the experiments such that the value function $V\in\mathbb{R}^{|{\cal S}|}$ lies in the span of the feature matrix. To achieve, we determine the optimal parameter $\theta^{\star}$ first, by sampling it entry-wise i.i.d. from $N(0,1)$. This produces the value function $V=\Phi\theta$, from which we reverse engineer the reward function
$R=(I-\gamma P^{\pi})V.$
The discount factor $\gamma$ is 0.5.
All algorithms start with zero initialization, i.e. $\theta_{0}=\omega_{0}=f_{0}=g_{0}=0$. The step sizes $\alpha_{k}$ and $\beta_{k}$ are set to constant values $45e-$ 4 and $32e-$ 3 for simplicity, while the step size $\lambda_{k}$ is selected as
$\lambda_{k}=\frac{4}{5(k+10)}.$